### v1

# DepthSplat: Connecting Gaussian Splatting and Depth

Abstract
Gaussian splatting and single/multi-view depth estimation are typically studied in isolation. In this paper, we present DepthSplat to connect Gaussian splatting and depth estimation and study their interactions. More specifically, we first contribute a robust multi-view depth model by leveraging pre-trained monocular depth features, leading to high-quality feed-forward 3D Gaussian splatting reconstructions. We also show that Gaussian splatting can serve as an unsupervised pre-training objective for learning powerful depth models from large-scale unlabelled datasets. We validate the synergy between Gaussian splatting and depth estimation through extensive ablation and cross-task transfer experiments. Our DepthSplat achieves state-of-the-art performance on ScanNet, RealEstate10K and DL3DV datasets in terms of both depth estimation and novel view synthesis, demonstrating the mutual benefits of connecting both tasks. Our code, models, and video results are available at https://haofeixu.github.io/depthsplat/.
Figure 1: DepthSplat enables cross-task interactions between Gaussian splatting and depth. Left: Better depth leads to improved Gaussian splatting reconstruction. Right: Unsupervised depth pre-training with Gaussian splatting leads to reduced depth prediction error.
1 Introduction
Novel view synthesis (Buehler et al., 2001; Zhou et al., 2018) and depth prediction (Schönberger et al., 2016; Eigen et al., 2014) are two fundamental tasks in computer vision, serving as the driving force behind numerous applications ranging from augmented reality to robotics and autonomous driving. There have been notable advancements in both areas recently.
For novel view synthesis, 3D Gaussian Splatting (3DGS) (Kerbl et al., 2023) has emerged as a popular technique due to its impressive real-time performance while attaining high visual fidelity. Recently, advances in feed-forward 3DGS models (Charatan et al., 2024; Chen et al., 2024; Szymanowicz et al., 2024) have been introduced to alleviate the need for tedious per-scene optimization, also enabling few-view 3D reconstruction. The state-of-the-art sparse-view method MVSplat (Chen et al., 2024) relies on feature matching-based multi-view depth estimation (Xu et al., 2023) to localize the 3D Gaussian positions, which makes it suffer from similar limitations (e.g., occlusions, texture-less regions, and reflective surfaces) as other multi-view depth methods (Schönberger et al., 2016; Yao et al., 2018; Gu et al., 2020; Wang et al., 2021; Duzceker et al., 2021).
On the other hand, significant progress has been made in monocular depth estimation, with recent models (Yang et al., 2024a; Ke et al., 2024; Yin et al., 2023; Fu et al., 2024; Ranftl et al., 2020; Eftekhar et al., 2021) achieving robust predictions on diverse in-the-wild data. However, these depths typically lack consistency across views, constraining their performance in downstream tasks (Wang et al., 2023; Yin et al., 2022). In addition, both state-of-the-art multi-view (Cao et al., 2022; Xu et al., 2023) and monocular (Yang et al., 2024a; Piccinelli et al., 2024; Yang et al., 2024b) depth models are trained with ground truth depth supervision, which prevents exploiting large unlabelled datasets for more robust depth predictions.
The integration of 3DGS with single/multi-view depth estimation presents a compelling solution to overcome the individual limitations of each technique while at the same time enhancing their strengths. To this end, we introduce DepthSplat, which exploits the complementary nature of sparse-view feed-forward 3DGS and robust single/multi-view depth estimation to improve the performance for both tasks.
Specifically, we first contribute a robust multi-view depth model by integrating pre-trained monocular depth features (Yang et al., 2024b) to the multi-view feature matching branch, which not only maintains the consistency of multi-view depth models but also leads to more robust results in situations that are hard to match (e.g., occlusions, texture-less regions and reflective surfaces). The predicted multi-view depth maps are then unprojected to 3D as the Gaussian centers, and we use an additional lightweight network to predict other remaining Gaussian parameters. They are combined together to achieve novel view synthesis with the splatting operation (Kerbl et al., 2023).
Thanks to our improved multi-view depth model, the quality of novel view synthesis with Gaussian splatting is also significantly enhanced (see Fig. 1 left). In addition, our Gaussian splatting module is fully differentiable, which requires only photometric supervision to optimize all model components. This provides a new, unsupervised way to pre-train depth prediction models on large-scale unlabeled datasets without requiring ground truth geometry information. The pre-trained depth model can be further fine-tuned for specific depth tasks and achieves superior results over training from scratch (see Fig. 1 right, where unsupervised pre-training leads to improved performance).
We conduct extensive experiments on the large-scale TartanAir (Wang et al., 2020), ScanNet (Dai et al., 2017) and RealEstate10K (Zhou et al., 2018) datasets for depth estimation and Gaussian splatting tasks, as well as the recently introduced DL3DV (Ling et al., 2023) dataset, which features complex real-world scenes and thus is more challenging. Under various evaluation settings, our DepthSplat achieves state-of-the-art results. The strong performance on both tasks demonstrates the mutual benefits of connecting Gaussian splatting and depth estimation.
2 Related Work
Multi-View Depth Estimation. As a core component of classical multi-view stereo pipelines (Schönberger et al., 2016), multi-view depth estimation exploits multi-view photometric consistency across multiple images to perform feature matching and predict the depth map of the reference image. Recently, many learning-based methods (Yao et al., 2018; Gu et al., 2020; Wang et al., 2021; Duzceker et al., 2021; Wang et al., 2022; Ding et al., 2022; Cao et al., 2022) have been proposed to improve depth accuracy. For example, MVSNet (Yao et al., 2018) uses the plane-sweep algorithm (Collins, 1996) to build a 3D cost volume, regularizes it with a 3D CNN, and then regresses the depth map. Though these learning-based methods significantly improve the depth quality when compared to traditional methods (Galliani et al., 2015; Schönberger et al., 2016; Xu & Tao, 2019), they cannot handle challenging situations where the multi-view photometric consistency assumption is not guaranteed, e.g., occlusions, low-textured areas, and non-Lambertian surfaces.
Monocular Depth Estimation. Recently, we have witnessed significant progress in depth estimation from a single image (Ranftl et al., 2020; Bhat et al., 2023; Yin et al., 2023; Yang et al., 2024a; Ke et al., 2024), and existing methods can produce surprisingly accurate results on diverse in-the-wild data. However, monocular depth methods inherently suffer from scale ambiguities and are not able to produce multi-view consistent depth predictions, which are crucial for downstream tasks like 3D reconstruction (Yin et al., 2022) and video depth estimation (Wang et al., 2023). In this paper, we leverage the powerful features from a pre-trained monocular depth model (Yang et al., 2024b) to augment feature-matching based multi-view depth estimation, which not only maintains high multi-view consistency but also leads to significantly improved robustness in challenging situations such as low-textured regions and reflective surfaces.
Monocular and Multi-View Depth Fusion. Several previous methods (Bae et al., 2022; Li et al., 2023; Yang et al., 2022; Cheng et al., 2024) try to fuse monocular and multi-view depths to alleviate the limitations of feature matching-based multi-view depth estimations. However, they either fuse single- and multi-view depth predictions with additional networks or rely on sophisticated architectures. In contrast, we identify the power of off-the-shell pre-trained monocular depth models and propose to augment multi-view cost volumes with monocular features, which not only leads to a simple model architecture but also strong performance.
Feed-Forward Gaussian Splatting. Several feed-forward 3D Gaussian splatting models (Charatan et al., 2024; Szymanowicz et al., 2024; Chen et al., 2024; Wewer et al., 2024; Tang et al., 2024; Xu et al., 2024b; Zhang et al., 2024) have been proposed in literature thanks to its efficiency and ability to handle sparse views. In particular, pixelSplat (Charatan et al., 2024) and Splatter Image (Szymanowicz et al., 2024) predict 3D Gaussians from image features, while MVSplat (Chen et al., 2024) encodes the feature matching information with cost volumes and achieves better geometry. However, it inherently suffers from the limitation of feature matching in challenging situations like texture-less regions and reflective surfaces. In this paper, we propose to integrate monocular features from pre-trained monocular depth models for more robust depth prediction and Gaussian splatting reconstruction. We additionally study the interactions between depth and Gaussian splatting tasks. Another line of work like LGM (Tang et al., 2024), GRM (Xu et al., 2024b), and GS-LRM (Zhang et al., 2024) relies significantly on the training data and compute, discarding explicit feature matching cues and learning everything purely from data. This makes them expensive to train (e.g., GS-LRM (Zhang et al., 2024) is trained with 64 A100 GPUs for 2 days), while our model can be trained in 1 day with 8 GPUs. Moreover, our Gaussian splatting module, in return, enables pre-training depth model from large-scale unlabelled datasets without the need for ground truth depths.
Depth and Gaussian Splatting. In addition to the experimental setup (feed-forward) studied in this paper, recently, there has been another line of work (Chung et al., 2024; Turkulainen et al., 2024) which applies an additional depth loss in the Gaussian splatting optimization process. We remark that these two approaches (feed-forward vs. per-scene optimization) are orthogonal. In particular, our approach focuses on exploring the advanced network architectures and the power of large-scale training data, while the optimization methods mainly study the effect of loss functions for regularizing the optimization process.
3 DepthSplat
Figure 2: DepthSplat connects depth estimation and 3D Gaussian splatting with a shared architecture, which enables cross-task transfer. In particular, DepthSplat consists of a multi-view branch to model feature-matching information and a single-view branch to extract monocular features. The per-view cost volumes and monocular features are concatenated for depth regression with a 2D U-Net architecture. For the depth estimation task, we train the depth model with ground truth depth supervision. For the Gaussian splatting task, we first unproject all depth maps to 3D as the Gaussian centers, and in parallel, we use an additional head to predict the remaining Gaussian parameters. Novel views are rendered with the splatting operation. The full model for novel view synthesis is trained with the photometric rendering loss, which can also be used as an unsupervised pre-training stage for the depth model.
Given $N$ input images $\{{\bm{I}}^{i}\}_{i=1}^{N},({\bm{I}}^{i}\in\mathbb{R}^{H\times W\times 3}$, where $H$ and $W$ are the image sizes) with corresponding projection matrices $\{\mathbf{P}_{i}\}_{i=1}^{N}$, ($\mathbf{P}_{i}\in\mathbb{R}^{3\times 4}$, computed from the intrinsic and extrinsic matrices), our goal is to predict dense per-pixel depth $\mathbf{D}_{i}\in\mathbb{R}^{H\times W}$ and per-pixel Gaussian parameters $\{(\bm{\mu}_{j},\alpha_{j},\bm{\Sigma}_{j},\bm{c}_{j})\}_{j=1}^{H\times W\times N}$ for each image, where $\bm{\mu}_{j}$, $\alpha_{j}$, $\bm{\Sigma}_{j}$ and $\bm{c}_{j}$ are the 3D Gaussian’s position, opacity, covariance, and color information. As shown in Fig. 2, the core of our method is a multi-view depth model augmented with monocular depth features, where we obtain the position $\bm{\mu}_{j}$ of each Gaussian by unprojecting depth to 3D with camera parameters, and other Gaussian parameters are predicted by an additional lightweight head.
More specifically, our depth model consists of two branches: one for modeling feature matching information using cost volumes, and another for extracting monocular features from a pre-trained monocular depth network. The cost volumes and monocular features are concatenated together for subsequent depth regression with a 2D U-Net and a softmax layer. For the depth task, we train our depth model with ground truth depth supervision. Our full model for novel view synthesis is trained with the photometric rendering loss, which can also be used as an unsupervised pre-training stage for the depth model. In the following, we introduce the individual components.
3.1 Multi-View Feature Matching
In this branch, we extract multi-view features with a multi-view Transformer architecture and then build multiple cost volumes that correspond to each input view.
Multi-View Feature Extraction. For $N$ input images, we first use a lightweight weight-sharing ResNet (He et al., 2016) architecture to get $s\times$ downsampled features for each image independently. To handle different image resolutions, we make the downsampling factor $s$ flexible by controlling the number of stride-2 $3\times 3$ convolutions. For example, the downsampling factor $s$ is 4 if two stride-2 convolutions are used and 8 if three are used. To exchange information across different views, we use a multi-view Swin Transformer (Liu et al., 2021; Xu et al., 2022; 2023) which contains six stacked self- and cross-attention layers to obtain multi-view-aware features $\{{\bm{F}_{i}}\}_{i=1}^{N}$, ${\bm{F}}^{i}\in\mathbb{R}^{\frac{H}{s}\times\frac{W}{s}\times C}$, where $C$ is the feature dimension. More specifically, cross-attention is performed between each reference view and other views. When more than two images ($N>2$) are given as input, we perform cross-attention between each reference view and its top-$2$ nearest neighboring views, which are selected based on their camera position distances to the reference view. This makes the computation tractable while maintaining cross-view interactions.
Feature Matching. We encode the feature matching information across different views with the plane-sweep stereo approach (Collins, 1996; Xu et al., 2023). More specifically, for each view $i$, we first uniformly sample $D$ depth candidates $\{d_{m}\}_{m=1}^{D}$ from the near and far depth ranges and then warp the feature ${\bm{F}}^{j}$ of view $j$ to view $i$ with the camera projection matrix and each depth candidate $d_{m}$. Then we obtain $D$ warped features $\{{\bm{F}}^{j\to i}_{d_{m}}\}_{m=1}^{D}$ that correspond to feature ${\bm{F}}^{i}$. We then measure their feature correlations with the dot-product operation (Xu & Zhang, 2020; Chen et al., 2024). The cost volume ${\bm{C}}_{i}\in\mathbb{R}^{\frac{H}{s}\times\frac{W}{s}\times D}$ for image $i$ is obtained by stacking all correlations. Accordingly, we obtain cost volumes $\{{\bm{C}}_{i}\}_{i=1}^{N}$ for all input images $\{{\bm{I}}_{i}\}_{i=1}^{N}$. For more than two input views, similar to the strategy in cross-view attention computation, we select the top-$2$ nearest views for each reference view and compute feature correlations with only the selected views. This enables our cost volume construction to achieve a good speed-accuracy trade-off and scale efficiently to a larger number of input views. The correlation values for the two selected views are combined with averaging.
3.2 Monocular Depth Feature Extraction
Despite the remarkable progress in multi-view feature matching-based depth estimation (Yao et al., 2018; Wang et al., 2022; Xu et al., 2023) and Gaussian splatting (Chen et al., 2024), they inherently suffer from limitations in challenging situations like occlusions, texture-less regions, and reflective surfaces. Thus, we propose to integrate pre-trained monocular depth features into the cost volume representation to handle scenarios that are challenging or impossible to match.
More specifically, we leverage the pre-trained monocular depth backbone from the recent Depth Anything (Yang et al., 2024b) model thanks to its impressive performance on diverse in-the-wild data. The monocular backbone is a ViT (Dosovitskiy et al., 2020; Oquab et al., 2023) model, which has a patch size of 14 and outputs a feature map that is $1/14$ spatial resolution of the original image. We simply bilinearly interpolate the spatial resolution of the monocular features to the same resolution as the cost volume in Sec. 3.1 and obtain the monocular feature ${\bm{F}}^{i}_{\mathrm{mono}}\in\mathbb{R}^{\frac{H}{s}\times\frac{W}{s}\times C_{\mathrm{mono}}}$ for input image ${\bm{I}}_{i}$, where $C_{\mathrm{mono}}$ is the dimension of the monocular feature. This process is performed for all input images in parallel and we obtain monocular features $\{{\bm{F}}^{i}_{\mathrm{mono}}\in\mathbb{R}^{\frac{H}{s}\times\frac{W}{s}\times C_{\mathrm{mono}}}\}_{i=1}^{N}$, which are subsequently used for per-view depth map estimations.
3.3 Feature Fusion and Depth Regression
To achieve robust and multi-view consistent depth predictions, we combine the monocular feature ${\bm{F}}^{i}_{\mathrm{mono}}\in\mathbb{R}^{\frac{H}{s}\times\frac{W}{s}\times C_{\mathrm{mono}}}$ and cost volume ${\bm{C}}_{i}\in\mathbb{R}^{\frac{H}{s}\times\frac{W}{s}\times D}$ via simple concatenation in the channel dimension. A subsequent 2D U-Net (Rombach et al., 2022; Ronneberger et al., 2015) is used to regress depth from the concatenated monocular features and cost volumes. This process is performed for all the input images in parallel and for each image, it outputs a tensor of shape $\frac{H}{s}\times\frac{W}{s}\times D$, where $D$ is the number of depth candidates. We then normalize the $D$ dimension with the softmax operation and perform a weighted average of all depth candidates to obtain the depth output.
We also apply a hierarchical matching (Gu et al., 2020) architecture where an additional refinement step at 2×2$\times$2 × higher feature resolution is employed to improve the performance further. More specifically, based on the coarse depth prediction, we perform a correspondence search on the 2×2$\times$2 × higher feature maps within the neighbors of the 2×2$\times$2 × upsampled coarse depth prediction. Since we already have a coarse depth prediction, we only need to search a smaller range at the higher resolution, and thus, we construct a smaller cost volume compared to the coarse stage. Such a $2$-scale hierarchical architecture not only leads to improved efficiency since most computation is spent on low resolution, but also leads to better results thanks to the use of higher-resolution features (Gu et al., 2020). Similar feature fusion and depth regression procedure is used to get higher-resolution depth predictions, which are subsequently upsampled to the full resolution with a learned upsampler (Ranftl et al., 2021).
3.4 Gaussian Parameter Prediction
For the task of 3D Gaussian splatting, we directly unproject the per-pixel depth maps to 3D with the camera parameters as the Gaussian centers $\bm{\mu}_{j}$. We append an additional lightweight network to predict other Gaussian parameters $\alpha_{j},\bm{\Sigma}_{j},\bm{c}_{j}$, which are opacity, covariance, and color, respectively. With all the predicted 3D Gaussians, we can render novel view images with the Gaussian splatting operation (Kerbl et al., 2023).
3.5 Training Loss
We study the properties of our proposed model on two tasks: depth estimation and novel view synthesis with 3D Gaussian splatting (Kerbl et al., 2023). The loss functions are introduced below.
Depth estimation. We train our depth model (without the Gaussian splatting head) with $\ell_{1}$ loss and gradient loss between the inverse depths of prediction and ground truth:
$L_{\mathrm{depth}}=\alpha\cdot|{\bm{D}}_{\mathrm{pred}}-{\bm{D}}_{\mathrm{gt}}|+\beta\cdot(|\partial_{x}{\bm{D}}_{\mathrm{pred}}-\partial_{x}{\bm{D}}_{\mathrm{gt}}|+|\partial_{y}{\bm{D}}_{\mathrm{pred}}-\partial_{y}{\bm{D}}_{\mathrm{gt}}|),$ (1)
where $\partial_{x}$ and $\partial_{y}$ denotes the gradients on the $x$ and $y$ directions, respectively. Following UniMatch (Xu et al., 2023), we use $\alpha=20$ and $\beta=20$.
View synthesis. We train our full model with a combination of mean squared error (MSE) and LPIPS (Zhang et al., 2018) losses between rendered and ground truth image colors:
$L_{\mathrm{gs}}=\sum_{m=1}^{M}\left(\mathrm{MSE}(I^{m}_{\mathrm{render}},I^{m}_{\mathrm{gt}})+\lambda\cdot\mathrm{LPIPS}(I^{m}_{\mathrm{render}},I^{m}_{\mathrm{gt}})\right),$ (2)
where $M$ is the number of novel views to render in a single forward pass. The LPIPS loss weight $\lambda$ is set to 0.05 following MVSplat (Chen et al., 2024).
4 Experiments
Implementation Details. We implement our method in PyTorch (Paszke et al., 2019) and optimize our model with the AdamW (Loshchilov & Hutter, 2017) optimizer and cosine learning rate learning rate schedule with warm up in the first $5\%$ of the total training iterations. We adopt the xFormers (Lefaudeux et al., 2022) library for our monocular ViT backbone implementation. We use a lower learning rate $2\times 10^{-6}$ for the pre-trained monocular backbone, and other remaining layers use a learning rate of $2\times 10^{-4}$. The feature downsampling factor $s$ in our multi-view branch (Sec. 3.1) is chosen based on the image resolution. More specifically, for experiments on the $256\times 256$ resolution RealEstate10K (Zhou et al., 2018) dataset, we choose $s=4$. For higher resolution datasets (e.g., TartanAir (Wang et al., 2020), ScanNet (Dai et al., 2017), KITTI (Geiger et al., 2013) and DL3DV (Ling et al., 2023)), we choose $s=8$. Our hierarchical matching models in Sec. 3.3 use 2-scale features, i.e., $1/8$ and $1/4$, or $1/4$ and $1/2$ resolutions.
Training Details. For depth experiments, we mainly follow the setup of UniMatch (Xu et al., 2023) for fair comparisons. More specifically, we train our depth model on 4 GH200 GPUs for 100K iterations with a total batch size of 32. For Gaussian splatting experiments on RealEstate10K (Zhou et al., 2018), we use the same training and testing splits of pixelSplat (Charatan et al., 2024) and MVSplat (Chen et al., 2024), and train our model on 8 GH200 GPUs for 150K iterations with a total batch size of 32, which takes about 1 day. For experiments on the DL3DV (Ling et al., 2023) dataset, we evaluate on the official benchmark split with 140 scenes, while other remaining 9896 scenes are used for training. We fine-tune our RealEstate10K pre-trained model on 4 A100 GPUs for 100K iterations with a total batch size of 4, where the number of input views is randomly sampled from 2 to 6. We evaluate the model’s performance on different number of input views (2, 4, 6). Our code and training scripts are available at https://github.com/cvg/depthsplat.
Table 1: DepthSplat model variants. We explore different monocular backbones and different multi-view models ($1$-scale and 2-scale features for hierarchical matching as described in Sec. 3.3), where both larger monocular backbones and $2$-scale hierarchical models lead to consistently improved performance for both depth estimation and view synthesis tasks.
Monocular Multi-View Depth (TartanAir) 3DGS (RealEstate10K) Param (M)
Abs Rel $\downarrow$ $\delta_{1}\uparrow$ PSNR $\uparrow$ SSIM $\uparrow$ LPIPS $\downarrow$
ViT-S $1$-scale 8.46 93.02 26.76 0.877 0.123 37
ViT-B $1$-scale 6.94 94.46 27.09 0.881 0.119 113
ViT-L $1$-scale 6.07 95.52 27.34 0.885 0.118 354
ViT-S $2$-scale 7.01 94.56 26.96 0.880 0.122 40
ViT-B $2$-scale 6.22 95.31 27.27 0.885 0.120 117
ViT-L $2$-scale 5.57 96.07 27.44 0.887 0.119 360
Table 2: Ablations. We evaluate the contribution of the monocular feature branch and the cost volume branch, as well as different monocular features. Our results indicate that the monocular feature and cost volume are complementary, with large performance drops when removing either one. The pre-trained Depth Anything network weights achieve overall the best performance.
Module Method Depth (TartanAir) 3DGS (RealEstate10K)
Abs Rel $\downarrow$ $\delta_{1}\uparrow$ PSNR $\uparrow$ SSIM $\uparrow$ LPIPS $\downarrow$
Components full 8.46 93.02 26.76 0.877 0.123
w/o mono feature 12.25 88.00 26.27 0.866 0.130
w/o cost volume 11.34 90.02 23.09 0.761 0.187
single branch 11.26 90.84 25.99 0.858 0.134
Monocular features ConvNeXt-T 10.50 91.13 26.28 0.867 0.130
Midas 9.53 91.61 26.40 0.869 0.129
DINO v2 8.93 92.49 26.68 0.874 0.125
Depth Anything v1 8.38 93.23 26.70 0.875 0.125
Depth Anything v2 8.46 93.02 26.76 0.877 0.123
Figure 3: Effect of monocular features for depth on ScanNet. The monocular feature greatly improves challenging situations like texture-less regions (e.g., the wall in the first example) and reflective surfaces (e.g., the refrigerator in the second example).
Figure 4: Effect of monocular features for 3DGS on RealEstate10K. Without monocular features, the model struggles at predicting reliable depth for pixels that are not able to find correspondences (e.g., the lounger chair highlighted with the read rectangle), which subsequently causes misalignment in the rendered image due to the incorrect geometry.
4.1 Model Variants
We first study several different model variants for both depth estimation and view synthesis tasks. In particular, we explore different monocular backbones (Yang et al., 2024b) (ViT-S, ViT-B, ViT-L) and different multi-view models ($1$-scale and $2$-scale). We conduct depth experiments on the large-scale TartanAir (Wang et al., 2020) synthetic dataset, which features both indoor and outdoor scenes and has perfect ground truth depth. The Gaussian splatting experiments are conducted on the standard RealEstate10K (Zhou et al., 2018) dataset. Following community standards, we report the depth evaluation metrics (Eigen et al., 2014) of Abs Rel (relative $\ell_{1}$ error) and $\delta_{1}$ (percentage of correctly estimated pixels within a threshold) and novel view synthesis metrics (Kerbl et al., 2023) of PSNR, SSIM, and LPIPS. The results in Table 1 demonstrate that both larger monocular backbones and $2$-scale hierarchical models lead to consistently improved performance for both tasks.
From Table 1, we can also see that better depth network architecture leads to improved view synthesis. In Appendix A.1, we conduct additional experiments to study the effect of different initializations for the depth network to the view synthesis performance. Our results show that a better depth model initialization also contributes to improved rendering quality. In summary, both better depth network architecture and better depth model initialization lead to improved novel view synthesis results.
4.2 Ablation Study and Analysis
In this section, we study the properties of our key components on the TartanAir dataset (for depth estimation) and RealEstate10K dataset (for view synthesis with Gaussian splatting).
Monocular Features for Depth Estimation and Gaussian Splatting. In Table 2, we compare our full model (full) with the model variant where the monocular depth feature branch (with a ViT-S model pre-trained by Depth Anything v2 (Yang et al., 2024b)) is removed (w/o mono feature), leaving only the multi-view branch. We can observe a clear performance drop for both depth and view synthesis tasks. In Fig. 3, we visualize the depth predictions and error maps of both models on the ScanNet dataset. The pure multi-view feature matching-based approach struggles a lot at texture-less regions and reflective surfaces. At the same time, our full model achieves reliable results thanks to the powerful prior captured in monocular depth features. We also show the visual comparisons for the Gaussian splatting task in Fig. 4 with two input views. For regions (e.g., the lounge chairs) that only appear in a single image, the pure multi-view method is unable to find correspondences. It thus produces unreliable depth predictions, leading to misalignment in the rendered novel views due to the incorrect geometry.
We also experiment with removing the cost volume (w/o cost volume) in the multi-view branch and observed a significant performance drop. This indicates that obtaining scale- and multi-view consistent predictions with a pure monocular depth backbone is challenging, which constrains achieving high-quality results in downstream tasks.
Fusion Strategy. We compare with alternative strategies for fusing the monocular features for multi-view depth estimation. In particular, we compare with MVSFormer (Cao et al., 2022) which constructs the cost volume with monocular features. More specifically, we replace our multi-view feature extractor with a weight-sharing ViT model and use the ViT features to build the cost volume as done in MVSFormer. This leads to a single-branch architecture (single branch in Table 2), unlike our two-branch design where the monocular features are not used to build the cost volume. We can observe that our fusion strategy performs significantly better than the single-branch design, potentially because our two-branch design disentangles feature matching and obtaining monocular priors, which makes the learning task easier.
Different Monocular Features. In Table 2, we also evaluate different monocular features, including the ConvNeXt-T (Liu et al., 2022) features used in AFNet (Cheng et al., 2024), and other popular monocular features including Midas (Ranftl et al., 2020) and DINO v2 (Oquab et al., 2023). The pre-trained Depth Anything monocular features achieve overall the best performance.
Table 3: Unsupervised Depth Pre-Training with Gaussian Splatting. We first perform unsupervised pre-training with Gaussian splatting on RealEstate10K and then measure the depth performance on TartanAir, ScanNet and KITTI after fine-tuning for the depth task. Compared to previous supervised pre-trained models Depth Anything (for monocular ViT) and UniMatch (for multi-view Transformer), our new unsupervised pre-training improves performance consistently in all metrics. The improvements are especially significant on the challenging datasets like TartanAir and KITTI.
Initialization TartanAir ScanNet KITTI
Abs Rel $\downarrow$ $\delta_{1}\uparrow$ Abs Rel $\downarrow$ $\delta_{1}\uparrow$ Abs Rel $\downarrow$ $\delta_{1}\uparrow$
Depth Anything (only mono) 11.12 89.97 6.78 96.13 11.67 85.96
UniMatch + Depth Anything (mv & mono) 10.86 90.55 6.70 96.14 11.56 87.27
DepthSplat (full model) 10.20 91.10 6.60 96.27 10.68 89.92
4.3 Unsupervised Depth Pre-Training with Gaussian Splatting
By connecting Gaussian splatting and depth, our DepthSplat provides a way to pre-train the depth model in a fully unsupervised manner. In particular, we first train our full model on the large-scale unlabelled RealEstate10K dataset (contains $\sim$ 67K Youtube videos) with only the Gaussian splatting rendering loss (Eqn. 2), without any direct supervision on the depth predictions. After pre-training, we take the pre-trained depth model and further fine-tune it to the depth task on the mixed TartanAir and VKITTI2 (Cabon et al., 2020) datasets with ground truth depth supervision.
In Table 3, we evaluate the performance on both in-domain TartanAir test set and the zero-shot generalization on unseen ScanNet and KITTI datasets. We compare with only initializing the monocular backbone with Depth Anything and additionally initializing the multi-view Transformer backbone with UniMatch (Xu et al., 2023). Our approach achieves the best results on all three datasets. It is also interesting to observe that our pre-training contributes more on more challenging datasets (i.e., TartanAir and KITTI which feature complex large-scale scenes unlike ScanNet only contains indoor scenes). We also note that both Depth Anything and UniMatch are trained with ground truth supervision, while our DepthSplat is trained with only photometric losses. Given the increasing popularity of view synthesis (Zheng & Vedaldi, 2024; Weng et al., 2023) and multi-view generative models (Shi et al., 2023; Blattmann et al., 2023), new multi-view datasets (Ling et al., 2023) and models (Voleti et al., 2024) are gradually introduced, our approach provides a way to pre-train depth models on large-scale unlabelled multi-view image datasets. This could potentially further improve the multi-view consistency and robustness of existing depth models (Yang et al., 2024a; Piccinelli et al., 2024), which are usually trained with ground truth depth supervision.
4.4 Benchmark Comparisons
Comparisons on ScanNet and RealEstate10K. Table 5 and Table 5 compare the depth and novel view synthesis results on the standard ScanNet and RealEstate10K benchmarks, respectively. For both comparisons, the numbers of input images are two. We can observe that our DepthSplat outperforms previous methods Teed & Deng (2019); Yu et al. (2021); Xu et al. (2024a); Chen et al. (2024) by clear margins on both datasets for both tasks. The visual comparison with previous best methods is shown in Fig. 5 and Fig. 6, where our method significantly improves the performance on challenging scenarios like texture-less regions and occlusions.
Comparisons on DL3DV. To further evaluate the performance on complex real-world scenes, we conduct comparisons with the latest state-of-the-art Gaussian splatting model MVSplat (Chen et al., 2024) on the recently introduced DL3DV dataset (Ling et al., 2023). We also compare the results of different numbers of input views (2, 4 and 6) on this dataset. We fine-tune MVSplat and our RealEstate10K pre-trained models on DL3DV training scenes and report the results on the benchmark test set in Table 6. Our DepthSplat consistently outperforms MVSplat in all metrics. Visual comparisons with MVSplat on the DL3DV dataset are shown in Fig. 7, where MVSplat’s depth quality lags behind our DepthSplat due to matching failure, which leads to blurry and distorted view synthesis results. We show more visual comparison results in Appendix A.2. It is also worth noting that our method scales more efficiently to more input views thanks to our lightweight local feature matching approach (Sec 3.1), which is unlike the expensive pair-wise matching used in MVSplat.
We invite the readers to our project page https://haofeixu.github.io/depthsplat/ for the video results on different number of input views (6 and 12), where our DepthSplat is able to reconstruct larger-scale or 360 scenes from more input views, without any optimization.
Table 4: Two-view Depth Estimation on ScanNet. Our DepthSplat outperforms all prior methods by clear margins.
Method Abs Rel $\downarrow$ RMSE $\downarrow$ RMSElog $\downarrow$
DeMoN 0.231 0.761 0.289
BA-Net 0.161 0.346 0.214
DeepV2D 0.057 0.168 0.080
NeuralRecon 0.047 0.164 0.093
DRO 0.053 0.168 0.081
UniMatch 0.059 0.179 0.082
DepthSplat (ours) 0.044 0.119 0.059
Table 5: Two-view 3DGS on RealEstate10K. Our DepthSplat achieves the best performance.
Method PSNR $\uparrow$ SSIM $\uparrow$ LPIPS $\downarrow$
pixelNeRF 20.43 0.589 0.550
GPNR 24.11 0.793 0.255
AttnRend 24.78 0.820 0.213
MuRF 26.10 0.858 0.143
pixelSplat 25.89 0.858 0.142
MVSplat 26.39 0.869 0.128
DepthSplat (ours) 27.44 0.887 0.119
Figure 5: Depth Comparison on ScanNet. Our DepthSplat performs significantly better than UniMatch (Xu et al., 2023) on challenging parts (e.g., edges of the couch and pillows).
Figure 6: Visual Synthesis on RealEstate10K. Our DepthSplat performs significantly better than pixelSplat (Charatan et al., 2024) and MVSplat (Chen et al., 2024) in challenging regions.
Table 6: Comparisons on DL3DV. Our DepthSplat not only consistently outperforms MVSplat on different number of input views, but also scales more efficiently to more input views.
Method #Views PSNR $\uparrow$ SSIM $\uparrow$ LPIPS $\downarrow$ Time (s)
MVSplat 2 17.54 0.529 0.402 0.072
DepthSplat 19.05 0.610 0.313 0.101
MVSplat 4 21.63 0.721 0.233 0.146
DepthSplat 22.82 0.766 0.188 0.124
MVSplat 6 22.93 0.775 0.193 0.263
DepthSplat 23.83 0.808 0.158 0.161
Figure 7: Visual Comparisons on DL3DV. Our DepthSplat performs significantly better than MVSplat (Chen et al., 2024) on regions that are hard to match (e.g., repeated patterns).
5 Conclusion
In this paper, we introduce DepthSplat, a new approach to connecting Gaussian splatting and depth to achieve state-of-the-art results on ScanNet, RealEstate10K and DL3DV datasets for both depth and view synthesis tasks. We also show that our model enables unsupervised pre-training depth with Gaussian splatting rendering loss, providing a way to leverage large-scale unlabelled multi-view image datasets for training more multi-view consistent and robust depth models. Our current model requires camera pose information as input along with the multi-view images, removing this requirement would be exciting future work.
Acknowledgement
We thank Yuedong Chen for his generous help with the DL3DV dataset. Andreas Geiger was supported by the ERC Starting Grant LEGO-3D (850533) and the DFG EXC number 2064/1 - project number 390727645.
Appendix A Appendix
A.1 Depth Pre-Training for Gaussian Splatting
We further study the effect of depth pre-training for Gaussian splatting experiments. Unlike the pre-trained models Depth Anything (Yang et al., 2024b) and UniMatch (Xu et al., 2023), which are trained for either monocular and multi-view features separately, we perform joint training of our full two-branch depth model on the depth datasets. We then compare the results of different initializations to the depth network for the full Gaussian splatting model. We can see from Table 7 that improved depth initialization leads to better view synthesis results.
Table 7: Depth to Gaussian Splatting Transfer. We compare different pre-trained network weights for initializing the depth network when training our full DepthSplat model for view synthesis. Compared to using Depth Anything and UniMatch pre-trained monocular and multi-view network weights for initializing the monocular ViT and multi-view Transformer features, our jointly trained 2-branch depth model (full model) performs best on all metrics.
Initialization PSNR $\uparrow$ SSIM $\uparrow$ LPIPS $\downarrow$
Depth Anything (only mono) 26.59 0.874 0.1256
UniMatch + Depth Anything (mv & mono) 26.76 0.877 0.1234
DepthSplat (full model) 26.81 0.878 0.1225
A.2 More Visual Comparisons on DL3DV
We show more visual comparison results on DL3DV in Fig. 8 with different number of input views, where our DepthSplat consistently outperforms MVSplat in challenging regions.
Figure 8: Different number of input views on DL3DV. Our DepthSplat performs consistently better than MVSplat (Chen et al., 2024).

### v2

@@
 More specifically, we first contribute a robust multi-view depth model by leveraging pre-trained monocular depth features, leading to high-quality feed-forward 3D Gaussian splatting reconstructions.
-We also show that Gaussian splatting can serve as an unsupervised pre-training objective for learning powerful depth models from large-scale unlabelled datasets.
+We also show that Gaussian splatting can serve as an unsupervised pre-training objective for learning powerful depth models from large-scale unlabeled datasets.
 We validate the synergy between Gaussian splatting and depth estimation through extensive ablation and cross-task transfer experiments.
 Our DepthSplat achieves state-of-the-art performance on ScanNet, RealEstate10K and DL3DV datasets in terms of both depth estimation and novel view synthesis, demonstrating the mutual benefits of connecting both tasks.
-Our code, models, and video results are available at https://haofeixu.github.io/depthsplat/.
 Figure 1: DepthSplat enables cross-task interactions between Gaussian splatting and depth.
@@
 1 Introduction
-Novel view synthesis (Buehler et al., 2001; Zhou et al., 2018) and depth prediction (Schönberger et al., 2016; Eigen et al., 2014) are two fundamental tasks in computer vision, serving as the driving force behind numerous applications ranging from augmented reality to robotics and autonomous driving.
+Novel view synthesis [4, 75] and depth prediction [41, 18] are two fundamental tasks in computer vision, serving as the driving force behind numerous applications ranging from augmented reality to robotics and autonomous driving.
 There have been notable advancements in both areas recently.
-For novel view synthesis, 3D Gaussian Splatting (3DGS) (Kerbl et al., 2023) has emerged as a popular technique due to its impressive real-time performance while attaining high visual fidelity.
-Recently, advances in feed-forward 3DGS models (Charatan et al., 2024; Chen et al., 2024; Szymanowicz et al., 2024) have been introduced to alleviate the need for tedious per-scene optimization, also enabling few-view 3D reconstruction.
-The state-of-the-art sparse-view method MVSplat (Chen et al., 2024) relies on feature matching-based multi-view depth estimation (Xu et al., 2023) to localize the 3D Gaussian positions, which makes it suffer from similar limitations (e.g., occlusions, texture-less regions, and reflective surfaces) as other multi-view depth methods (Schönberger et al., 2016; Yao et al., 2018; Gu et al., 2020; Wang et al., 2021; Duzceker et al., 2021).
-On the other hand, significant progress has been made in monocular depth estimation, with recent models (Yang et al., 2024a; Ke et al., 2024; Yin et al., 2023; Fu et al., 2024; Ranftl et al., 2020; Eftekhar et al., 2021) achieving robust predictions on diverse in-the-wild data.
-However, these depths typically lack consistency across views, constraining their performance in downstream tasks (Wang et al., 2023; Yin et al., 2022).
-In addition, both state-of-the-art multi-view (Cao et al., 2022; Xu et al., 2023) and monocular (Yang et al., 2024a; Piccinelli et al., 2024; Yang et al., 2024b) depth models are trained with ground truth depth supervision, which prevents exploiting large unlabelled datasets for more robust depth predictions.
+For novel view synthesis, 3D Gaussian Splatting (3DGS) [27] has emerged as a popular technique due to its impressive real-time performance while attaining high visual fidelity.
+Recently, advances in feed-forward 3DGS models [7, 8, 46] have been introduced to alleviate the need for tedious per-scene optimization, also enabling few-view 3D reconstruction.
+The state-of-the-art sparse-view method MVSplat [8] relies on feature matching-based multi-view depth estimation [61] to localize the 3D Gaussian positions, which makes it suffer from similar limitations (e.g., occlusions, texture-less regions, and reflective surfaces) as other multi-view depth methods [41, 67, 23, 53, 16].
+On the other hand, significant progress has been made in monocular depth estimation, with recent models [65, 26, 69, 20, 37, 17] achieving robust predictions on diverse in-the-wild data.
+However, these depths typically lack consistent scales across multiple views, constraining their performance in downstream tasks like 3D reconstruction [56, 68].
 The integration of 3DGS with single/multi-view depth estimation presents a compelling solution to overcome the individual limitations of each technique while at the same time enhancing their strengths.
-To this end, we introduce DepthSplat, which exploits the complementary nature of sparse-view feed-forward 3DGS and robust single/multi-view depth estimation to improve the performance for both tasks.
-Specifically, we first contribute a robust multi-view depth model by integrating pre-trained monocular depth features (Yang et al., 2024b) to the multi-view feature matching branch, which not only maintains the consistency of multi-view depth models but also leads to more robust results in situations that are hard to match (e.g., occlusions, texture-less regions and reflective surfaces).
+To this end, we introduce DepthSplat, which exploits the complementary nature of sparse-view feed-forward 3DGS and robust monocular depth estimation to improve the performance for both tasks.
+Specifically, we first contribute a robust multi-view depth model by integrating pre-trained monocular depth features [66] to the multi-view feature matching branch, which not only maintains the consistency of multi-view depth models but also leads to more robust results in situations that are hard to match (e.g., occlusions, texture-less regions and reflective surfaces).
 The predicted multi-view depth maps are then unprojected to 3D as the Gaussian centers, and we use an additional lightweight network to predict other remaining Gaussian parameters.
-They are combined together to achieve novel view synthesis with the splatting operation (Kerbl et al., 2023).
+They are combined together to achieve novel view synthesis with the splatting operation [27].
+While previous methods [1, 29, 9] also try to fuse monocular and multi-view depths, they usually rely on sophisticated architectures.
+In contrast, we identify the power of off-the-shelf pre-trained monocular depth models and propose to augment multi-view cost volumes with monocular features, leading to a simpler model and stronger performance.
 Thanks to our improved multi-view depth model, the quality of novel view synthesis with Gaussian splatting is also significantly enhanced (see Fig.
@@
 1 right, where unsupervised pre-training leads to improved performance).
-We conduct extensive experiments on the large-scale TartanAir (Wang et al., 2020), ScanNet (Dai et al., 2017) and RealEstate10K (Zhou et al., 2018) datasets for depth estimation and Gaussian splatting tasks, as well as the recently introduced DL3DV (Ling et al., 2023) dataset, which features complex real-world scenes and thus is more challenging.
+We conduct extensive experiments on the large-scale TartanAir [55], ScanNet [12] and RealEstate10K [75] datasets for depth estimation and Gaussian splatting tasks, as well as the recently introduced DL3DV [30] dataset, which features complex real-world scenes and thus is more challenging.
 Under various evaluation settings, our DepthSplat achieves state-of-the-art results.
-The strong performance on both tasks demonstrates the mutual benefits of connecting Gaussian splatting and depth estimation.
+We also achieve promising feed-forward reconstruction results of large-scale or 360 scenes from up to 12 input views at $512\times 960$ resolutions.
+The strong performance on both tasks demonstrates the mutual benefits of connecting Gaussian splatting and depth.
 2 Related Work
 Multi-View Depth Estimation.
-As a core component of classical multi-view stereo pipelines (Schönberger et al., 2016), multi-view depth estimation exploits multi-view photometric consistency across multiple images to perform feature matching and predict the depth map of the reference image.
-Recently, many learning-based methods (Yao et al., 2018; Gu et al., 2020; Wang et al., 2021; Duzceker et al., 2021; Wang et al., 2022; Ding et al., 2022; Cao et al., 2022) have been proposed to improve depth accuracy.
-For example, MVSNet (Yao et al., 2018) uses the plane-sweep algorithm (Collins, 1996) to build a 3D cost volume, regularizes it with a 3D CNN, and then regresses the depth map.
-Though these learning-based methods significantly improve the depth quality when compared to traditional methods (Galliani et al., 2015; Schönberger et al., 2016; Xu & Tao, 2019), they cannot handle challenging situations where the multi-view photometric consistency assumption is not guaranteed, e.g., occlusions, low-textured areas, and non-Lambertian surfaces.
+As a core component of classical multi-view stereo pipelines [41], multi-view depth estimation exploits multi-view photometric consistency across multiple images to perform feature matching and predict the depth map of the reference image.
+In recent years, many learning-based methods [67, 23, 53, 16, 54, 13, 6] have been proposed to improve depth accuracy.
+Though these learning-based methods significantly improve the depth quality compared to traditional methods [21, 41, 63], they cannot handle challenging situations where the multi-view photometric consistency assumption is not guaranteed, e.g., occlusions, low-textured areas, and non-Lambertian surfaces.
 Monocular Depth Estimation.
-Recently, we have witnessed significant progress in depth estimation from a single image (Ranftl et al., 2020; Bhat et al., 2023; Yin et al., 2023; Yang et al., 2024a; Ke et al., 2024), and existing methods can produce surprisingly accurate results on diverse in-the-wild data.
-However, monocular depth methods inherently suffer from scale ambiguities and are not able to produce multi-view consistent depth predictions, which are crucial for downstream tasks like 3D reconstruction (Yin et al., 2022) and video depth estimation (Wang et al., 2023).
-In this paper, we leverage the powerful features from a pre-trained monocular depth model (Yang et al., 2024b) to augment feature-matching based multi-view depth estimation, which not only maintains high multi-view consistency but also leads to significantly improved robustness in challenging situations such as low-textured regions and reflective surfaces.
-Monocular and Multi-View Depth Fusion.
-Several previous methods (Bae et al., 2022; Li et al., 2023; Yang et al., 2022; Cheng et al., 2024) try to fuse monocular and multi-view depths to alleviate the limitations of feature matching-based multi-view depth estimations.
-However, they either fuse single- and multi-view depth predictions with additional networks or rely on sophisticated architectures.
-In contrast, we identify the power of off-the-shell pre-trained monocular depth models and propose to augment multi-view cost volumes with monocular features, which not only leads to a simple model architecture but also strong performance.
-Feed-Forward Gaussian Splatting.
-Several feed-forward 3D Gaussian splatting models (Charatan et al., 2024; Szymanowicz et al., 2024; Chen et al., 2024; Wewer et al., 2024; Tang et al., 2024; Xu et al., 2024b; Zhang et al., 2024) have been proposed in literature thanks to its efficiency and ability to handle sparse views.
-In particular, pixelSplat (Charatan et al., 2024) and Splatter Image (Szymanowicz et al., 2024) predict 3D Gaussians from image features, while MVSplat (Chen et al., 2024) encodes the feature matching information with cost volumes and achieves better geometry.
-However, it inherently suffers from the limitation of feature matching in challenging situations like texture-less regions and reflective surfaces.
-In this paper, we propose to integrate monocular features from pre-trained monocular depth models for more robust depth prediction and Gaussian splatting reconstruction.
-We additionally study the interactions between depth and Gaussian splatting tasks.
-Another line of work like LGM (Tang et al., 2024), GRM (Xu et al., 2024b), and GS-LRM (Zhang et al., 2024) relies significantly on the training data and compute, discarding explicit feature matching cues and learning everything purely from data.
-This makes them expensive to train (e.g., GS-LRM (Zhang et al., 2024) is trained with 64 A100 GPUs for 2 days), while our model can be trained in 1 day with 8 GPUs.
-Moreover, our Gaussian splatting module, in return, enables pre-training depth model from large-scale unlabelled datasets without the need for ground truth depths.
-Depth and Gaussian Splatting.
-In addition to the experimental setup (feed-forward) studied in this paper, recently, there has been another line of work (Chung et al., 2024; Turkulainen et al., 2024) which applies an additional depth loss in the Gaussian splatting optimization process.
-We remark that these two approaches (feed-forward vs.
-per-scene optimization) are orthogonal.
-In particular, our approach focuses on exploring the advanced network architectures and the power of large-scale training data, while the optimization methods mainly study the effect of loss functions for regularizing the optimization process.
-3 DepthSplat
+Recently, we have witnessed significant progress in depth estimation from a single image [37, 2, 69, 65, 26], and existing methods can produce surprisingly accurate results on diverse in-the-wild data.
+However, monocular depth methods inherently suffer from scale ambiguities, which makes it challenging to use them for downstream tasks like 3D reconstruction [68].
+In this paper, we leverage the powerful features from a pre-trained monocular depth model [66] to augment feature-matching based multi-view depth estimation, which not only maintains the multi-view consistency but also leads to significantly improved robustness in challenging situations such as low-textured regions and reflective surfaces.
 Figure 2: DepthSplat connects depth estimation and 3D Gaussian splatting with a shared architecture, which enables cross-task transfer.
@@
 The full model for novel view synthesis is trained with the photometric rendering loss, which can also be used as an unsupervised pre-training stage for the depth model.
+Feed-Forward Gaussian Splatting.
+Several feed-forward 3D Gaussian splatting models [7, 46, 8, 58, 48, 64, 72] have been proposed in literature thanks to its efficiency and ability to handle sparse views.
+In particular, pixelSplat [7] and Splatter Image [46] predict 3D Gaussians from image features, while MVSplat [8] encodes the feature matching information with cost volumes and achieves better geometry.
+However, it inherently suffers from the limitation of feature matching in challenging situations like texture-less regions and reflective surfaces.
+In this paper, we propose to integrate monocular features from pre-trained monocular depth models [66] for more robust depth prediction and 3D Gaussian reconstruction.
+Another line of work like LGM [48], GRM [64], and GS-LRM [72] relies significantly on the training data and compute, discarding explicit feature matching cues and learning priors purely from data.
+This makes them expensive to train (e.g., GS-LRM [72] is trained with 64 A100 GPUs for 2 days), while our model can be trained in 20 hours with 8 GPUs.
+Moreover, our Gaussian splatting module, in return, enables pre-training depth model from large-scale unlabeled datasets without the need for ground truth depths.
+Depth and Gaussian Splatting.
+Flash3D [45] explores a pre-trained monocular depth model [36] for single-image 3D Gaussian reconstruction.
+In contrast, we target large-scale scene reconstruction from multiple views.
+TranSplat [71] leverages monocular depth to improve Gaussian reconstruction.
+However, a crucial difference is that TranSplat uses monocular features to refine the coarse depth predicted from a cost volume, which makes it inherently vulnerable to the error in the coarse stage and, thus, suffers from the error propagation issue.
+In contrast, we avoid this by combining cost volume and monocular features early, which also enables our method to perform significantly better than TranSplat (Tab.
+5).
+There has been another line of work [10, 50] which applies an additional depth loss in the Gaussian splatting optimization process.
+We note that these two approaches (feed-forward vs.
+per-scene optimization) are orthogonal.
+3 DepthSplat
 Given $N$ input images $\{{\bm{I}}^{i}\}_{i=1}^{N},({\bm{I}}^{i}\in\mathbb{R}^{H\times W\times 3}$, where $H$ and $W$ are the image sizes) with corresponding projection matrices $\{\mathbf{P}_{i}\}_{i=1}^{N}$, ($\mathbf{P}_{i}\in\mathbb{R}^{3\times 4}$, computed from the intrinsic and extrinsic matrices), our goal is to predict dense per-pixel depth $\mathbf{D}_{i}\in\mathbb{R}^{H\times W}$ and per-pixel Gaussian parameters $\{(\bm{\mu}_{j},\alpha_{j},\bm{\Sigma}_{j},\bm{c}_{j})\}_{j=1}^{H\times W\times N}$ for each image, where $\bm{\mu}_{j}$, $\alpha_{j}$, $\bm{\Sigma}_{j}$ and $\bm{c}_{j}$ are the 3D Gaussian’s position, opacity, covariance, and color information.
@@
 Multi-View Feature Extraction.
-For $N$ input images, we first use a lightweight weight-sharing ResNet (He et al., 2016) architecture to get $s\times$ downsampled features for each image independently.
+For $N$ input images, we first use a lightweight weight-sharing ResNet [25] architecture to get $s\times$ downsampled features for each image independently.
 To handle different image resolutions, we make the downsampling factor $s$ flexible by controlling the number of stride-2 $3\times 3$ convolutions.
 For example, the downsampling factor $s$ is 4 if two stride-2 convolutions are used and 8 if three are used.
-To exchange information across different views, we use a multi-view Swin Transformer (Liu et al., 2021; Xu et al., 2022; 2023) which contains six stacked self- and cross-attention layers to obtain multi-view-aware features $\{{\bm{F}_{i}}\}_{i=1}^{N}$, ${\bm{F}}^{i}\in\mathbb{R}^{\frac{H}{s}\times\frac{W}{s}\times C}$, where $C$ is the feature dimension.
+To exchange information across different views, we use a multi-view Swin Transformer [31, 60, 61] which contains six stacked self- and cross-attention layers to obtain multi-view-aware features $\{{\bm{F}_{i}}\}_{i=1}^{N}$, ${\bm{F}}^{i}\in\mathbb{R}^{\frac{H}{s}\times\frac{W}{s}\times C}$, where $C$ is the feature dimension.
 More specifically, cross-attention is performed between each reference view and other views.
-When more than two images ($N>2$) are given as input, we perform cross-attention between each reference view and its top-$2$ nearest neighboring views, which are selected based on their camera position distances to the reference view.
+When more than two images ($N>2$) are given as input, we perform cross-attention between each reference view and its top-$2$ nearest neighboring views, which are selected based on their camera distances to the reference view.
 This makes the computation tractable while maintaining cross-view interactions.
 Feature Matching.
-We encode the feature matching information across different views with the plane-sweep stereo approach (Collins, 1996; Xu et al., 2023).
+We encode the feature matching information across different views with the plane-sweep stereo approach [11, 61].
 More specifically, for each view $i$, we first uniformly sample $D$ depth candidates $\{d_{m}\}_{m=1}^{D}$ from the near and far depth ranges and then warp the feature ${\bm{F}}^{j}$ of view $j$ to view $i$ with the camera projection matrix and each depth candidate $d_{m}$.
 Then we obtain $D$ warped features $\{{\bm{F}}^{j\to i}_{d_{m}}\}_{m=1}^{D}$ that correspond to feature ${\bm{F}}^{i}$.
-We then measure their feature correlations with the dot-product operation (Xu & Zhang, 2020; Chen et al., 2024).
+We then measure their feature correlations with the dot-product operation [59, 8].
 The cost volume ${\bm{C}}_{i}\in\mathbb{R}^{\frac{H}{s}\times\frac{W}{s}\times D}$ for image $i$ is obtained by stacking all correlations.
@@
 3.2 Monocular Depth Feature Extraction
-Despite the remarkable progress in multi-view feature matching-based depth estimation (Yao et al., 2018; Wang et al., 2022; Xu et al., 2023) and Gaussian splatting (Chen et al., 2024), they inherently suffer from limitations in challenging situations like occlusions, texture-less regions, and reflective surfaces.
-Thus, we propose to integrate pre-trained monocular depth features into the cost volume representation to handle scenarios that are challenging or impossible to match.
-More specifically, we leverage the pre-trained monocular depth backbone from the recent Depth Anything (Yang et al., 2024b) model thanks to its impressive performance on diverse in-the-wild data.
-The monocular backbone is a ViT (Dosovitskiy et al., 2020; Oquab et al., 2023) model, which has a patch size of 14 and outputs a feature map that is $1/14$ spatial resolution of the original image.
+Despite the remarkable progress in multi-view feature matching-based depth estimation [67, 54, 61] and Gaussian splatting [8], they inherently suffer from limitations in challenging situations like occlusions, texture-less regions, and reflective surfaces.
+Thus, we propose to integrate pre-trained monocular depth features into the cost volume to handle scenarios that are challenging or impossible to match.
+More specifically, we leverage the pre-trained monocular depth backbone from the recent Depth Anything V2 [66] model thanks to its impressive performance on diverse in-the-wild data.
+The monocular backbone is a ViT [14, 34] model, which has a patch size of 14 and outputs a feature map that is $1/14$ spatial resolution of the original image.
 We simply bilinearly interpolate the spatial resolution of the monocular features to the same resolution as the cost volume in Sec.
@@
 To achieve robust and multi-view consistent depth predictions, we combine the monocular feature ${\bm{F}}^{i}_{\mathrm{mono}}\in\mathbb{R}^{\frac{H}{s}\times\frac{W}{s}\times C_{\mathrm{mono}}}$ and cost volume ${\bm{C}}_{i}\in\mathbb{R}^{\frac{H}{s}\times\frac{W}{s}\times D}$ via simple concatenation in the channel dimension.
-A subsequent 2D U-Net (Rombach et al., 2022; Ronneberger et al., 2015) is used to regress depth from the concatenated monocular features and cost volumes.
+A subsequent 2D U-Net [39, 40] is used to regress depth from the concatenated monocular features and cost volumes.
 This process is performed for all the input images in parallel and for each image, it outputs a tensor of shape $\frac{H}{s}\times\frac{W}{s}\times D$, where $D$ is the number of depth candidates.
-We then normalize the $D$ dimension with the softmax operation and perform a weighted average of all depth candidates to obtain the depth output.
-We also apply a hierarchical matching (Gu et al., 2020) architecture where an additional refinement step at 2×2$\times$2 × higher feature resolution is employed to improve the performance further.
+We normalize the $D$ dimension with the softmax operation and perform a weighted average of all depth candidates to obtain the depth output.
+We also apply a hierarchical matching [23] architecture where an additional refinement step at 2×2$\times$2 × higher feature resolution is employed to improve the performance further.
 More specifically, based on the coarse depth prediction, we perform a correspondence search on the 2×2$\times$2 × higher feature maps within the neighbors of the 2×2$\times$2 × upsampled coarse depth prediction.
 Since we already have a coarse depth prediction, we only need to search a smaller range at the higher resolution, and thus, we construct a smaller cost volume compared to the coarse stage.
-Such a $2$-scale hierarchical architecture not only leads to improved efficiency since most computation is spent on low resolution, but also leads to better results thanks to the use of higher-resolution features (Gu et al., 2020).
-Similar feature fusion and depth regression procedure is used to get higher-resolution depth predictions, which are subsequently upsampled to the full resolution with a learned upsampler (Ranftl et al., 2021).
+Such a $2$-scale hierarchical architecture not only leads to improved efficiency since most computation is spent on low resolution, but also leads to better results thanks to the use of higher-resolution features [23].
+Similar feature fusion and depth regression procedure is used to get higher-resolution depth predictions, which are subsequently upsampled [38] to the full resolution.
 3.4 Gaussian Parameter Prediction
 For the task of 3D Gaussian splatting, we directly unproject the per-pixel depth maps to 3D with the camera parameters as the Gaussian centers $\bm{\mu}_{j}$.
-We append an additional lightweight network to predict other Gaussian parameters $\alpha_{j},\bm{\Sigma}_{j},\bm{c}_{j}$, which are opacity, covariance, and color, respectively.
-With all the predicted 3D Gaussians, we can render novel view images with the Gaussian splatting operation (Kerbl et al., 2023).
+To predict other remaining Gaussian parameters $\alpha_{j}$ (opacity), $\bm{\Sigma}_{j}$ (covariance) and $\bm{c}_{j}$ (color), we use an additional lightweight 2D U-Net [39], where it takes the concatenated image, depth and feature information as input and outputs all the remaining Gaussian parameters.
+With all the predicted 3D Gaussians, we can render novel views with the Gaussian splatting operation [27].
 3.5 Training Loss
-We study the properties of our proposed model on two tasks: depth estimation and novel view synthesis with 3D Gaussian splatting (Kerbl et al., 2023). The loss functions are introduced below.
+We study the properties of our proposed model on two tasks: depth estimation and novel view synthesis with 3D Gaussian splatting [27]. The loss functions are introduced below.
 Depth estimation. We train our depth model (without the Gaussian splatting head) with $\ell_{1}$ loss and gradient loss between the inverse depths of prediction and ground truth:
-$L_{\mathrm{depth}}=\alpha\cdot|{\bm{D}}_{\mathrm{pred}}-{\bm{D}}_{\mathrm{gt}}|+\beta\cdot(|\partial_{x}{\bm{D}}_{\mathrm{pred}}-\partial_{x}{\bm{D}}_{\mathrm{gt}}|+|\partial_{y}{\bm{D}}_{\mathrm{pred}}-\partial_{y}{\bm{D}}_{\mathrm{gt}}|),$ (1)
-where $\partial_{x}$ and $\partial_{y}$ denotes the gradients on the $x$ and $y$ directions, respectively. Following UniMatch (Xu et al., 2023), we use $\alpha=20$ and $\beta=20$.
-View synthesis. We train our full model with a combination of mean squared error (MSE) and LPIPS (Zhang et al., 2018) losses between rendered and ground truth image colors:
+$L_{\mathrm{depth}}$ $=\alpha\cdot|{\bm{D}}_{\mathrm{pred}}-{\bm{D}}_{\mathrm{gt}}|+\beta\cdot|\partial_{x}{\bm{D}}_{\mathrm{pred}}-\partial_{x}{\bm{D}}_{\mathrm{gt}}|$
+$\quad+\beta\cdot|\partial_{y}{\bm{D}}_{\mathrm{pred}}-\partial_{y}{\bm{D}}_{\mathrm{gt}}|),$ (1)
+where $\partial_{x}$ and $\partial_{y}$ denotes the gradients on the $x$ and $y$ directions, respectively. Following UniMatch [61], we use $\alpha=20$ and $\beta=20$.
+View synthesis. We train our full model with a combination of mean squared error (MSE) and LPIPS [73] losses between rendered and ground truth image colors:
 $L_{\mathrm{gs}}=\sum_{m=1}^{M}\left(\mathrm{MSE}(I^{m}_{\mathrm{render}},I^{m}_{\mathrm{gt}})+\lambda\cdot\mathrm{LPIPS}(I^{m}_{\mathrm{render}},I^{m}_{\mathrm{gt}})\right),$ (2)
-where $M$ is the number of novel views to render in a single forward pass. The LPIPS loss weight $\lambda$ is set to 0.05 following MVSplat (Chen et al., 2024).
+where $M$ is the number of novel views to render in a single forward pass. The LPIPS loss weight $\lambda$ is set to 0.05 [8].
 4 Experiments
 Implementation Details.
-We implement our method in PyTorch (Paszke et al., 2019) and optimize our model with the AdamW (Loshchilov & Hutter, 2017) optimizer and cosine learning rate learning rate schedule with warm up in the first $5\%$ of the total training iterations.
-We adopt the xFormers (Lefaudeux et al., 2022) library for our monocular ViT backbone implementation.
-We use a lower learning rate $2\times 10^{-6}$ for the pre-trained monocular backbone, and other remaining layers use a learning rate of $2\times 10^{-4}$.
+We implement our method in PyTorch [35] and optimize our model with the AdamW [33] optimizer and cosine learning rate schedule.
+We adopt the xFormers [28] library for our monocular ViT backbone implementation.
+We use a lower learning rate $2\times 10^{-6}$ for the pre-trained Depth Anything V2 [66] backbone, and other remaining layers use a learning rate of $2\times 10^{-4}$.
 The feature downsampling factor $s$ in our multi-view branch (Sec.
 3.1) is chosen based on the image resolution.
-More specifically, for experiments on the $256\times 256$ resolution RealEstate10K (Zhou et al., 2018) dataset, we choose $s=4$.
-For higher resolution datasets (e.g., TartanAir (Wang et al., 2020), ScanNet (Dai et al., 2017), KITTI (Geiger et al., 2013) and DL3DV (Ling et al., 2023)), we choose $s=8$.
+More specifically, for experiments on the $256\times 256$ resolution RealEstate10K [75] dataset, we choose $s=4$.
+For higher resolution datasets (e.g., TartanAir [55], ScanNet [12], KITTI [22] and DL3DV [30]), we choose $s=8$.
 Our hierarchical matching models in Sec.
@@
 Training Details.
-For depth experiments, we mainly follow the setup of UniMatch (Xu et al., 2023) for fair comparisons.
+For depth experiments, we mainly follow the setup of UniMatch [61] for fair comparisons on the ScanNet [12] dataset.
 More specifically, we train our depth model on 4 GH200 GPUs for 100K iterations with a total batch size of 32.
-For Gaussian splatting experiments on RealEstate10K (Zhou et al., 2018), we use the same training and testing splits of pixelSplat (Charatan et al., 2024) and MVSplat (Chen et al., 2024), and train our model on 8 GH200 GPUs for 150K iterations with a total batch size of 32, which takes about 1 day.
-For experiments on the DL3DV (Ling et al., 2023) dataset, we evaluate on the official benchmark split with 140 scenes, while other remaining 9896 scenes are used for training.
+For the ablation experiments on depth task, we mainly use synthetic datasets (TartanAir [55] and VKITTI2 [5]) for training since they provide high-quality ground truth depths.
+For Gaussian splatting experiments, we consider both low-resolution and high-resolution images.
+In particular, for comparisons with previous methods [7, 8], we mainly use $256\times 256$ resolution RealEstate10K [75] and $256\times 448$ resolution DL3DV [30] datasets.
+We also report high-resolution ($512\times 960$) results on RealEstate10K and DL3DV datasets for qualitative evaluations.
+For experiments on the $256\times 256$ RealEstate10K dataset, we train our model on 8 GH200 GPUs for 150K iterations with a total batch size of 32, which takes about 20 hours.
+For experiments on the $256\times 448$ DL3DV [30] dataset, we evaluate on the official benchmark split with 140 scenes, while other remaining scenes are used for training.
 We fine-tune our RealEstate10K pre-trained model on 4 A100 GPUs for 100K iterations with a total batch size of 4, where the number of input views is randomly sampled from 2 to 6.
 We evaluate the model’s performance on different number of input views (2, 4, 6).
-Our code and training scripts are available at https://github.com/cvg/depthsplat.
+For high-resolution results, we fine-tune our low-resolution pre-trained models on high-resolution images.
+More training details are provided in the appendix.
+Our code and training scripts are available at github.com/cvg/depthsplat to ease reproducibility.
 Table 1: DepthSplat model variants.
-We explore different monocular backbones and different multi-view models ($1$-scale and 2-scale features for hierarchical matching as described in Sec.
-3.3), where both larger monocular backbones and $2$-scale hierarchical models lead to consistently improved performance for both depth estimation and view synthesis tasks.
-Monocular Multi-View Depth (TartanAir) 3DGS (RealEstate10K) Param (M)
+We evaluate different monocular backbones and different multi-view models ($1$-scale and 2-scale features for hierarchical matching as described in Sec.
+3.3), where both larger monocular backbones and $2$-scale hierarchical models lead to consistently improved performance for both depth and view synthesis tasks.
+Our model parameters, inference time, and memory consumption are provided in Tab.
+A.1 of the appendix.
+Monocular Multi-View Depth (TartanAir) 3DGS (RealEstate10K)
 Abs Rel $\downarrow$ $\delta_{1}\uparrow$ PSNR $\uparrow$ SSIM $\uparrow$ LPIPS $\downarrow$
-ViT-S $1$-scale 8.46 93.02 26.76 0.877 0.123 37
-ViT-B $1$-scale 6.94 94.46 27.09 0.881 0.119 113
-ViT-L $1$-scale 6.07 95.52 27.34 0.885 0.118 354
-ViT-S $2$-scale 7.01 94.56 26.96 0.880 0.122 40
-ViT-B $2$-scale 6.22 95.31 27.27 0.885 0.120 117
-ViT-L $2$-scale 5.57 96.07 27.44 0.887 0.119 360
+ViT-S $1$-scale 8.46 93.02 26.76 0.877 0.123
+ViT-B $1$-scale 6.94 94.46 27.09 0.881 0.119
+ViT-L $1$-scale 6.07 95.52 27.34 0.885 0.118
+ViT-S $2$-scale 7.01 94.56 26.96 0.880 0.122
+ViT-B $2$-scale 6.22 95.31 27.27 0.885 0.120
+ViT-L $2$-scale 5.57 96.07 27.44 0.887 0.119
 Table 2: Ablations.
-We evaluate the contribution of the monocular feature branch and the cost volume branch, as well as different monocular features.
+We evaluate the contribution of the monocular feature branch and the cost volume branch (1st group of experiments), as well as different monocular features (2nd group of experiments).
 Our results indicate that the monocular feature and cost volume are complementary, with large performance drops when removing either one.
-The pre-trained Depth Anything network weights achieve overall the best performance.
-Module Method Depth (TartanAir) 3DGS (RealEstate10K)
+The pre-trained Depth Anything V2 model weights achieve the best view synthesis results.
+Components Depth (TartanAir) 3DGS (RealEstate10K)
 Abs Rel $\downarrow$ $\delta_{1}\uparrow$ PSNR $\uparrow$ SSIM $\uparrow$ LPIPS $\downarrow$
-Components full 8.46 93.02 26.76 0.877 0.123
+full 8.46 93.02 26.76 0.877 0.123
 w/o mono feature 12.25 88.00 26.27 0.866 0.130
@@
 single branch 11.26 90.84 25.99 0.858 0.134
-Monocular features ConvNeXt-T 10.50 91.13 26.28 0.867 0.130
+ConvNeXt-T 10.50 91.13 26.28 0.867 0.130
 Midas 9.53 91.61 26.40 0.869 0.129
@@
 We first study several different model variants for both depth estimation and view synthesis tasks.
-In particular, we explore different monocular backbones (Yang et al., 2024b) (ViT-S, ViT-B, ViT-L) and different multi-view models ($1$-scale and $2$-scale).
-We conduct depth experiments on the large-scale TartanAir (Wang et al., 2020) synthetic dataset, which features both indoor and outdoor scenes and has perfect ground truth depth.
-The Gaussian splatting experiments are conducted on the standard RealEstate10K (Zhou et al., 2018) dataset.
-Following community standards, we report the depth evaluation metrics (Eigen et al., 2014) of Abs Rel (relative $\ell_{1}$ error) and $\delta_{1}$ (percentage of correctly estimated pixels within a threshold) and novel view synthesis metrics (Kerbl et al., 2023) of PSNR, SSIM, and LPIPS.
-The results in Table 1 demonstrate that both larger monocular backbones and $2$-scale hierarchical models lead to consistently improved performance for both tasks.
-From Table 1, we can also see that better depth network architecture leads to improved view synthesis.
-In Appendix A.1, we conduct additional experiments to study the effect of different initializations for the depth network to the view synthesis performance.
+In particular, we explore different monocular backbones [66] (ViT-S, ViT-B, ViT-L) and different multi-view models ($1$-scale and $2$-scale).
+We conduct depth experiments on the large-scale TartanAir [55] synthetic dataset, which features both indoor and outdoor scenes and has perfect ground truth depth.
+The Gaussian splatting experiments are conducted on the standard RealEstate10K [75] dataset.
+Following community standards, we report the depth evaluation metrics [18] of Abs Rel (relative $\ell_{1}$ error) and $\delta_{1}$ (percentage of correctly estimated pixels within a threshold) and novel view synthesis metrics [27] of PSNR, SSIM, and LPIPS.
+The results in Tab.
+1 demonstrate that both larger monocular backbones and $2$-scale hierarchical models lead to consistently improved performance for both tasks.
+From Tab.
+1, we can also observe that better depth network architecture leads to improved view synthesis.
+In Tab.
+B.1 of the appendix, we conduct additional experiments to study the effect of different initializations for the depth network to the view synthesis performance.
 Our results show that a better depth model initialization also contributes to improved rendering quality.
 In summary, both better depth network architecture and better depth model initialization lead to improved novel view synthesis results.
-4.2 Ablation Study and Analysis
-In this section, we study the properties of our key components on the TartanAir dataset (for depth estimation) and RealEstate10K dataset (for view synthesis with Gaussian splatting).
-Monocular Features for Depth Estimation and Gaussian Splatting.
-In Table 2, we compare our full model (full) with the model variant where the monocular depth feature branch (with a ViT-S model pre-trained by Depth Anything v2 (Yang et al., 2024b)) is removed (w/o mono feature), leaving only the multi-view branch.
+4.2 Ablation and Analysis
+In this section, we study the properties of our key components on the TartanAir dataset (for depth) and RealEstate10K dataset (for view synthesis with Gaussian splatting).
+Monocular Features for Depth and Gaussian Splatting.
+In Tab.
+2, we compare our full model (full) with the model variant where the monocular depth feature branch (with a ViT-S model pre-trained by Depth Anything V2 [66]) is removed (w/o mono feature), leaving only the multi-view branch.
 We can observe a clear performance drop for both depth and view synthesis tasks.
@@
 We also experiment with removing the cost volume (w/o cost volume) in the multi-view branch and observed a significant performance drop.
-This indicates that obtaining scale- and multi-view consistent predictions with a pure monocular depth backbone is challenging, which constrains achieving high-quality results in downstream tasks.
+This indicates that obtaining scale- and multi-view consistent predictions with a pure monocular depth backbone is challenging, which constrains achieving high-quality results for 3D Gaussian reconstruction.
 Fusion Strategy.
 We compare with alternative strategies for fusing the monocular features for multi-view depth estimation.
-In particular, we compare with MVSFormer (Cao et al., 2022) which constructs the cost volume with monocular features.
+In particular, we compare with MVSFormer [6] which constructs the cost volume with monocular features.
 More specifically, we replace our multi-view feature extractor with a weight-sharing ViT model and use the ViT features to build the cost volume as done in MVSFormer.
-This leads to a single-branch architecture (single branch in Table 2), unlike our two-branch design where the monocular features are not used to build the cost volume.
+This leads to a single-branch architecture (single branch in Tab.
+2), unlike our two-branch design where the monocular features are not used to build the cost volume.
 We can observe that our fusion strategy performs significantly better than the single-branch design, potentially because our two-branch design disentangles feature matching and obtaining monocular priors, which makes the learning task easier.
+Table 3: Unsupervised Depth Pre-Training with Gaussian Splatting.
+We evaluate the same architecture but with different weights.
+The pre-training is performed on RealEstate10K, and fine-tuning is performed on TartanAir and VKITTI2.
+Unsupervised pre-training (2nd row) produces improved depth predictions compared to random initialization (1st).
+The performance can be further improved with additional fine-tuning (4th), where the benefit of pre-training is especially significant on challenging datasets like TartanAir and KITTI compared to training from random initialization (3rd).
+Pre-Train Fine-Tune TartanAir ScanNet KITTI
+Abs Rel $\downarrow$ $\delta_{1}\uparrow$ Abs Rel $\downarrow$ $\delta_{1}\uparrow$ Abs Rel $\downarrow$ $\delta_{1}\uparrow$
+76.22 3.43 39.97 22.56 90.69 0.00
+29.53 53.16 21.51 76.09 56.83 46.26
+10.86 90.55 6.70 96.14 11.56 87.27
+10.20 91.10 6.60 96.27 10.68 89.92
+Table 4: Two-view Depth Estimation on ScanNet. Our DepthSplat outperforms all prior methods by clear margins.
+Method Abs Rel $\downarrow$ RMSE $\downarrow$ RMSElog $\downarrow$
+DeMoN [51] 0.231 0.761 0.289
+BA-Net [47] 0.161 0.346 0.214
+DeepV2D [49] 0.057 0.168 0.080
+NeuralRecon [44] 0.047 0.164 0.093
+DRO [24] 0.053 0.168 0.081
+UniMatch [61] 0.059 0.179 0.082
+DepthSplat 0.044 0.119 0.059
 Different Monocular Features.
-In Table 2, we also evaluate different monocular features, including the ConvNeXt-T (Liu et al., 2022) features used in AFNet (Cheng et al., 2024), and other popular monocular features including Midas (Ranftl et al., 2020) and DINO v2 (Oquab et al., 2023).
-The pre-trained Depth Anything monocular features achieve overall the best performance.
-Table 3: Unsupervised Depth Pre-Training with Gaussian Splatting.
-We first perform unsupervised pre-training with Gaussian splatting on RealEstate10K and then measure the depth performance on TartanAir, ScanNet and KITTI after fine-tuning for the depth task.
-Compared to previous supervised pre-trained models Depth Anything (for monocular ViT) and UniMatch (for multi-view Transformer), our new unsupervised pre-training improves performance consistently in all metrics.
-The improvements are especially significant on the challenging datasets like TartanAir and KITTI.
-Initialization TartanAir ScanNet KITTI
-Abs Rel $\downarrow$ $\delta_{1}\uparrow$ Abs Rel $\downarrow$ $\delta_{1}\uparrow$ Abs Rel $\downarrow$ $\delta_{1}\uparrow$
-Depth Anything (only mono) 11.12 89.97 6.78 96.13 11.67 85.96
-UniMatch + Depth Anything (mv & mono) 10.86 90.55 6.70 96.14 11.56 87.27
-DepthSplat (full model) 10.20 91.10 6.60 96.27 10.68 89.92
+In Tab.
+2, we also evaluate different monocular features, including the ConvNeXt-T [32] features used in AFNet [9], and other popular monocular features including Midas [37] and DINOv2 [34].
+The pre-trained Depth Anything V2 [66] monocular features achieve the best view synthesis results.
 4.3 Unsupervised Depth Pre-Training with Gaussian Splatting
 By connecting Gaussian splatting and depth, our DepthSplat provides a way to pre-train the depth model in a fully unsupervised manner.
-In particular, we first train our full model on the large-scale unlabelled RealEstate10K dataset (contains $\sim$ 67K Youtube videos) with only the Gaussian splatting rendering loss (Eqn.
+In particular, we first train our full model on the large-scale unlabeled RealEstate10K dataset (contains $\sim$ 67K Youtube videos) with only the Gaussian splatting rendering loss (Eq.
 2), without any direct supervision on the depth predictions.
-After pre-training, we take the pre-trained depth model and further fine-tune it to the depth task on the mixed TartanAir and VKITTI2 (Cabon et al., 2020) datasets with ground truth depth supervision.
-In Table 3, we evaluate the performance on both in-domain TartanAir test set and the zero-shot generalization on unseen ScanNet and KITTI datasets.
-We compare with only initializing the monocular backbone with Depth Anything and additionally initializing the multi-view Transformer backbone with UniMatch (Xu et al., 2023).
-Our approach achieves the best results on all three datasets.
-It is also interesting to observe that our pre-training contributes more on more challenging datasets (i.e., TartanAir and KITTI which feature complex large-scale scenes unlike ScanNet only contains indoor scenes).
-We also note that both Depth Anything and UniMatch are trained with ground truth supervision, while our DepthSplat is trained with only photometric losses.
-Given the increasing popularity of view synthesis (Zheng & Vedaldi, 2024; Weng et al., 2023) and multi-view generative models (Shi et al., 2023; Blattmann et al., 2023), new multi-view datasets (Ling et al., 2023) and models (Voleti et al., 2024) are gradually introduced, our approach provides a way to pre-train depth models on large-scale unlabelled multi-view image datasets.
-This could potentially further improve the multi-view consistency and robustness of existing depth models (Yang et al., 2024a; Piccinelli et al., 2024), which are usually trained with ground truth depth supervision.
+After pre-training, we take the pre-trained depth model and further fine-tune it to the depth task on the mixed TartanAir [55] and VKITTI2 [5] datasets with ground truth depth supervision.
+In Tab.
+3, we evaluate the performance on both in-domain TartanAir test set and the zero-shot generalization on unseen ScanNet and KITTI datasets.
+We can observe that the benefit of pre-training is especially significant on the challenging datasets like TartanAir and KITTI.
+The results align well with the observation in [19] which suggests that the pre-training acts as regularization and guides the learning towards better minima and enables better generalization to out-of-distribution datasets.
+The visual comparison results are provided in Fig.
+C.1 of the appendix.
+Given the increasing popularity of view synthesis [74, 57] and multi-view generative models [42, 3], new multi-view datasets [30] and models [52] are gradually introduced, our approach provides a way to pre-train depth models on large-scale unlabeled multi-view image datasets.
+Table 5: Two-view 3DGS on RealEstate10K. Our DepthSplat achieves the best performance.
+Method PSNR $\uparrow$ SSIM $\uparrow$ LPIPS $\downarrow$
+pixelNeRF [70] 20.43 0.589 0.550
+GPNR [43] 24.11 0.793 0.255
+AttnRend [15] 24.78 0.820 0.213
+MuRF [62] 26.10 0.858 0.143
+pixelSplat [7] 25.89 0.858 0.142
+MVSplat [8] 26.39 0.869 0.128
+TranSplat [71] 26.69 0.875 0.125
+DepthSplat 27.44 0.887 0.119
+Figure 5: Visual synthesis on RealEstate10K. Our DepthSplat performs significantly better than pixelSplat [7] and MVSplat [8] in challenging regions.
+Figure 6: Depth results on unseen samples. Our depth model generalizes robustly to different scene types.
 4.4 Benchmark Comparisons
 Comparisons on ScanNet and RealEstate10K.
-Table 5 and Table 5 compare the depth and novel view synthesis results on the standard ScanNet and RealEstate10K benchmarks, respectively.
-For both comparisons, the numbers of input images are two.
-We can observe that our DepthSplat outperforms previous methods Teed & Deng (2019); Yu et al.
-(2021); Xu et al.
-(2024a); Chen et al.
-(2024) by clear margins on both datasets for both tasks.
-The visual comparison with previous best methods is shown in Fig.
-5 and Fig.
-6, where our method significantly improves the performance on challenging scenarios like texture-less regions and occlusions.
+Tab.
+4 and Tab.
+5 compare the depth and novel view synthesis results on the standard ScanNet and RealEstate10K benchmarks, respectively.
+We can see clearly that our DepthSplat achieves state-of-the-art performance on both datasets for both tasks.
+The visual comparison with previous methods is shown in Fig.
+5, where our method significantly improves the performance on challenging scenarios like texture-less regions and occlusions.
+Table 6: Comparisons on DL3DV. Our DepthSplat not only consistently outperforms MVSplat on different number of input views, but also scales more efficiently to more input views.
+Method #Views PSNR $\uparrow$ SSIM $\uparrow$ LPIPS $\downarrow$ Time (s)
+MVSplat [8] 2 17.54 0.529 0.402 0.072
+DepthSplat 19.05 0.610 0.313 0.101
+MVSplat [8] 4 21.63 0.721 0.233 0.146
+DepthSplat 22.82 0.766 0.188 0.124
+MVSplat [8] 6 22.93 0.775 0.193 0.263
+DepthSplat 23.83 0.808 0.158 0.161
+Figure 7: Visual Comparisons on DL3DV. Our DepthSplat performs significantly better than MVSplat [8] on regions that hard to match (e.g., repeated patterns).
 Comparisons on DL3DV.
-To further evaluate the performance on complex real-world scenes, we conduct comparisons with the latest state-of-the-art Gaussian splatting model MVSplat (Chen et al., 2024) on the recently introduced DL3DV dataset (Ling et al., 2023).
+To evaluate the performance on complex real-world scenes, we conduct comparisons with the representative model MVSplat [8] on the recently introduced DL3DV dataset [30].
 We also compare the results of different numbers of input views (2, 4 and 6) on this dataset.
-We fine-tune MVSplat and our RealEstate10K pre-trained models on DL3DV training scenes and report the results on the benchmark test set in Table 6.
-Our DepthSplat consistently outperforms MVSplat in all metrics.
+We fine-tune MVSplat and our RealEstate10K pre-trained models on DL3DV training scenes and report the results on the benchmark test set in Tab.
+6.
+Our DepthSplat consistently outperforms MVSplat in all metrics, and our improvements are more significant with fewer input views, which indicates that our model is more robust to sparse views.
 Visual comparisons with MVSplat on the DL3DV dataset are shown in Fig.
 7, where MVSplat’s depth quality lags behind our DepthSplat due to matching failure, which leads to blurry and distorted view synthesis results.
-We show more visual comparison results in Appendix A.2.
-It is also worth noting that our method scales more efficiently to more input views thanks to our lightweight local feature matching approach (Sec 3.1), which is unlike the expensive pair-wise matching used in MVSplat.
-We invite the readers to our project page https://haofeixu.github.io/depthsplat/ for the video results on different number of input views (6 and 12), where our DepthSplat is able to reconstruct larger-scale or 360 scenes from more input views, without any optimization.
-Table 4: Two-view Depth Estimation on ScanNet. Our DepthSplat outperforms all prior methods by clear margins.
-Method Abs Rel $\downarrow$ RMSE $\downarrow$ RMSElog $\downarrow$
-DeMoN 0.231 0.761 0.289
-BA-Net 0.161 0.346 0.214
-DeepV2D 0.057 0.168 0.080
-NeuralRecon 0.047 0.164 0.093
-DRO 0.053 0.168 0.081
-UniMatch 0.059 0.179 0.082
-DepthSplat (ours) 0.044 0.119 0.059
-Table 5: Two-view 3DGS on RealEstate10K. Our DepthSplat achieves the best performance.
-Method PSNR $\uparrow$ SSIM $\uparrow$ LPIPS $\downarrow$
-pixelNeRF 20.43 0.589 0.550
-GPNR 24.11 0.793 0.255
-AttnRend 24.78 0.820 0.213
-MuRF 26.10 0.858 0.143
-pixelSplat 25.89 0.858 0.142
-MVSplat 26.39 0.869 0.128
-DepthSplat (ours) 27.44 0.887 0.119
-Figure 5: Depth Comparison on ScanNet. Our DepthSplat performs significantly better than UniMatch (Xu et al., 2023) on challenging parts (e.g., edges of the couch and pillows).
-Figure 6: Visual Synthesis on RealEstate10K. Our DepthSplat performs significantly better than pixelSplat (Charatan et al., 2024) and MVSplat (Chen et al., 2024) in challenging regions.
-Table 6: Comparisons on DL3DV. Our DepthSplat not only consistently outperforms MVSplat on different number of input views, but also scales more efficiently to more input views.
-Method #Views PSNR $\uparrow$ SSIM $\uparrow$ LPIPS $\downarrow$ Time (s)
-MVSplat 2 17.54 0.529 0.402 0.072
-DepthSplat 19.05 0.610 0.313 0.101
-MVSplat 4 21.63 0.721 0.233 0.146
-DepthSplat 22.82 0.766 0.188 0.124
-MVSplat 6 22.93 0.775 0.193 0.263
-DepthSplat 23.83 0.808 0.158 0.161
-Figure 7: Visual Comparisons on DL3DV. Our DepthSplat performs significantly better than MVSplat (Chen et al., 2024) on regions that are hard to match (e.g., repeated patterns).
+We show more visual comparison results in Fig.
+C.3 of the appendix.
+It is also worth noting that our method scales more efficiently to more input views thanks to our lightweight local feature matching approach (Sec.
+3.1), which is unlike the expensive global pair-wise matching used in MVSplat.
+Table 7: Cross-dataset generalization. When generalizing to DL3DV with RealEstate10K pre-trained models, our DepthSplat is consistently better than MVSplat for both small and large baselines.
+Method Small Baseline Large Baseline
+PSNR $\uparrow$ SSIM $\uparrow$ LPIPS $\downarrow$ PSNR $\uparrow$ SSIM $\uparrow$ LPIPS $\downarrow$
+MVSplat [8] 19.19 0.608 0.304 16.26 0.459 0.438
+DepthSplat 20.29 0.672 0.259 17.50 0.536 0.376
+Cross-dataset generalization.
+We further evaluate the cross-dataset generalization capability of our model by directly testing on DL3DV with RealEstate10K pre-trained models.
+The results are shown in Tab.
+7, where our DepthSplat consistently outperforms MVSplat for both small and large baseline inputs (visual comparisons are shown in Fig.
+C.2 of the appendix).
+We also show some visual results of our depth model on unseen samples in Fig.
+6, where our model generalizes robustly to different scene types.
+High-resolution results.
+In addition to the comparison results on the $256\times 256$ resolution RealEstate10K and $256\times 448$ DL3DV dataets, we show more high-resolution ($512\times 960$) qualitative results in Fig.
+C.4 and Fig.
+C.5 of the appendix.
+We also invite the readers to our project page haofeixu.github.io/depthsplat for high-resolution ($512\times 960$) video results on different number of input views (6 and 12), where our DepthSplat is able to reconstruct larger-scale or 360 scenes from more input views.
 5 Conclusion
 In this paper, we introduce DepthSplat, a new approach to connecting Gaussian splatting and depth to achieve state-of-the-art results on ScanNet, RealEstate10K and DL3DV datasets for both depth and view synthesis tasks.
-We also show that our model enables unsupervised pre-training depth with Gaussian splatting rendering loss, providing a way to leverage large-scale unlabelled multi-view image datasets for training more multi-view consistent and robust depth models.
+We also show that our model enables unsupervised pre-training depth with Gaussian splatting rendering loss, providing a way to leverage large-scale unlabeled multi-view image datasets for training more robust and generalizable multi-view depth models.
 Our current model requires camera pose information as input along with the multi-view images, removing this requirement would be exciting future work.
-Acknowledgement
+6 Acknowledgement
 We thank Yuedong Chen for his generous help with the DL3DV dataset. Andreas Geiger was supported by the ERC Starting Grant LEGO-3D (850533) and the DFG EXC number 2064/1 - project number 390727645.
-Appendix A Appendix
-A.1 Depth Pre-Training for Gaussian Splatting
-We further study the effect of depth pre-training for Gaussian splatting experiments.
-Unlike the pre-trained models Depth Anything (Yang et al., 2024b) and UniMatch (Xu et al., 2023), which are trained for either monocular and multi-view features separately, we perform joint training of our full two-branch depth model on the depth datasets.
-We then compare the results of different initializations to the depth network for the full Gaussian splatting model.
-We can see from Table 7 that improved depth initialization leads to better view synthesis results.
-Table 7: Depth to Gaussian Splatting Transfer.
-We compare different pre-trained network weights for initializing the depth network when training our full DepthSplat model for view synthesis.
-Compared to using Depth Anything and UniMatch pre-trained monocular and multi-view network weights for initializing the monocular ViT and multi-view Transformer features, our jointly trained 2-branch depth model (full model) performs best on all metrics.
+Appendix
+A Model Efficiency
+We report the number of parameters, inference time and memory consumption in Tab.
+A.1 for different model sizes.
+In particular, the “small” model refers to “ViT-S monocular, 1-scale multi-view” in Tab.
+1 of the main paper, and the “base” model refers to “ViT-B monocular, 2-scale multi-view”, and the “large” model refers to “ViT-L monocular, 2-scale multi-view”.
+We report the inference time (measured on an A100 GPU) and memory for 2 input views at $256\times 256$ and $512\times 512$ resolutions.
+Our DepthSplat can support different speed-accuracy trade-offs.
+Table A.1: Model efficiency.
+Model Size Resolution Params (M) Time (s) Memory (GB)
+Small $256\times 256$ 37 0.065 1.4
+Base 117 0.081 2.1
+Large 360 0.105 3.5
+Small $512\times 512$ 37 0.091 3.1
+Base 117 0.155 5.4
+Large 360 0.313 8.4
+B Depth Pre-Training for Gaussian Splatting
+In Sec.
+4.1 of the main paper, we have shown that better depth architectures lead to improved view synthesis with Gaussian splatting.
+We further study the effect of different weight initializations for the depth model.
+Specifically, we compare three variants: 1) only initializing the monocular feature with Depth Anything V2 [66]; 2) initializing the monocular feature with Depth Anything V2 [66] and the multi-view feature with UniMatch [61]; 3) initializing the full depth model by pre-training the depth model on TartanAir [55].
+We can observe from Tab.
+B.1 that better depth initialization also leads to improved view synthesis results.
+Table B.1: Depth pre-training for Gaussian splatting.
+We compare different weight initializations for the depth model when training our full DepthSplat model for view synthesis.
+Compared to 1) only initializing the monocular feature (mono features) and 2) initializing the monocular and multi-view features (mono & mv features), our pre-trained full depth model (full depth model) achieves the best view synthesis results.
 Initialization PSNR $\uparrow$ SSIM $\uparrow$ LPIPS $\downarrow$
-Depth Anything (only mono) 26.59 0.874 0.1256
-UniMatch + Depth Anything (mv & mono) 26.76 0.877 0.1234
-DepthSplat (full model) 26.81 0.878 0.1225
-A.2 More Visual Comparisons on DL3DV
-We show more visual comparison results on DL3DV in Fig. 8 with different number of input views, where our DepthSplat consistently outperforms MVSplat in challenging regions.
-Figure 8: Different number of input views on DL3DV. Our DepthSplat performs consistently better than MVSplat (Chen et al., 2024).
+mono features 26.59 0.874 0.1256
+mono & mv features 26.76 0.877 0.1234
+full depth model 26.81 0.878 0.1225
+C More Visual Results
+C.1 Unsupervised Depth Pre-Training with Gaussian Splatting
+In Fig.
+C.1, we show the comparisons of depth estimation results with and without Gaussian splatting pre-training.
+We observe that pre-training leads to better results in texture-less regions.
+We hypothesize that pre-training provides regularizations (as also observed in  [19]) to the challenging scenarios and accordingly leads to improved performance.
+Figure C.1: Effect of unsupervised depth pre-training. We observe that the unsupervised pre-training leads to improved performance for texture-less regions.
+C.2 Cross-Dataset Generalization
+In Fig.
+C.2, we show the cross-dataset generalization results on the DL3DV [30] dataset, which are obtained with the RealEstate10K [75] trained models.
+Our DepthSplat generalizes more robustly than MVSplat [8] on unseen scenes.
+Figure C.2: Generalization from RealEstate10K to DL3DV.
+C.3 Visual Comparisons on DL3DV
+In Fig. C.3, we compare the visual synthesis results from 4 input views with MVSplat [8]. Our DepthSplat better preserves the scene structures.
+Figure C.3: View synthesis from 4 input views on DL3DV.
+C.4 High-Resolution Results
+In Fig.
+C.4 and Fig.
+C.5, we show the view synthesis and multi-view depth estimation results at $512\times 960$ resolutions.
+Note that MVSplat [8] runs out-of-memory on such high resolutions with 6 or 12 input views, while our DepthSplat significantly improves the efficiency with two technical components.
+First, our lowest feature resolution is $1/8$ of the image resolution, while MVSplat uses $1/4$.
+Second, we use local cross-view attentions (detailed in Sec.
+3.1 of the main paper) unlike the pair-wise global attentions in MVSplat.
+Figure C.4: View Synthesis at $512\times 960$ resolutions from 6 input views.
+Figure C.5: Depth predictions on DL3DV with 12 input views. The image resolutions are $512\times 960$.
+D More Implementation Details
+We provide more implementation details on high-resolution experiments.
+For high-resolution experiments, we choose our small model which contains a ViT-S monocular branch and a single-scale multi-view branch.
+We first train our model on RealEstate10K at $448\times 768$ resolutions for 200K iterations with a batch size of 8, and then we fine-tune the model on DL3DV at $448\times 768$ resolutions for 200K iterations with a batch size of 8.
+We use random number of input views (2 to 6) for training.
+Despite that our model is trained at $448\times 768$ resolutions with at most 6 input views, we observe our model can generalize reasonably well on higher resolutions (e.g., $512\times 960$) and more input views (e.g., 12), as we show in our project page: haofeixu.github.io/depthsplat.

### v3

@@
 Abstract
-Gaussian splatting and single/multi-view depth estimation are typically studied in isolation.
+Gaussian splatting and single-view depth estimation are typically studied in isolation.
 In this paper, we present DepthSplat to connect Gaussian splatting and depth estimation and study their interactions.
 More specifically, we first contribute a robust multi-view depth model by leveraging pre-trained monocular depth features, leading to high-quality feed-forward 3D Gaussian splatting reconstructions.
-We also show that Gaussian splatting can serve as an unsupervised pre-training objective for learning powerful depth models from large-scale unlabeled datasets.
+We also show that Gaussian splatting can serve as an unsupervised pre-training objective for learning powerful depth models from large-scale multi-view posed datasets.
 We validate the synergy between Gaussian splatting and depth estimation through extensive ablation and cross-task transfer experiments.
 Our DepthSplat achieves state-of-the-art performance on ScanNet, RealEstate10K and DL3DV datasets in terms of both depth estimation and novel view synthesis, demonstrating the mutual benefits of connecting both tasks.
+In addition, DepthSplat enables feed-forward reconstruction from 12 input views ($512\times 960$ resolutions) in 0.6 seconds.
 Figure 1: DepthSplat enables cross-task interactions between Gaussian splatting and depth.
@@
 However, these depths typically lack consistent scales across multiple views, constraining their performance in downstream tasks like 3D reconstruction [59, 73].
-The integration of 3DGS with single/multi-view depth estimation presents a compelling solution to overcome the individual limitations of each technique while at the same time enhancing their strengths.
+The integration of 3DGS with single-view depth estimation presents a compelling solution to overcome the individual limitations of each technique while at the same time enhancing their strengths.
 To this end, we introduce DepthSplat, which exploits the complementary nature of sparse-view feed-forward 3DGS and robust monocular depth estimation to improve the performance for both tasks.
@@
 In addition, our Gaussian splatting module is fully differentiable, which requires only photometric supervision to optimize all model components.
-This provides a new, unsupervised way to pre-train depth prediction models on large-scale unlabeled datasets without requiring ground truth geometry information.
+This provides a new, unsupervised way to pre-train depth prediction models on large-scale multi-view posed datasets without requiring ground truth geometry information.
 The pre-trained depth model can be further fine-tuned for specific depth tasks and achieves superior results over training from scratch (see Fig.
@@
 Under various evaluation settings, our DepthSplat achieves state-of-the-art results.
-We also achieve promising feed-forward reconstruction results of large-scale or 360 scenes from up to 12 input views at $512\times 960$ resolutions.
+We also achieve promising feed-forward reconstruction results of large-scale or $360^{\circ}$ scenes from 12 input views ($512\times 960$ resolutions) in 0.6 seconds.
 The strong performance on both tasks demonstrates the mutual benefits of connecting Gaussian splatting and depth.
@@
 Feed-Forward Gaussian Splatting.
-Several feed-forward 3D Gaussian splatting models [7, 46, 8, 58, 48, 64, 72] have been proposed in literature thanks to its efficiency and ability to handle sparse views.
+Several feed-forward 3D Gaussian splatting models [7, 48, 9, 62, 51, 68, 77, 10, 8] have been proposed in literature thanks to its efficiency and ability to handle sparse views.
 In particular, pixelSplat [7] and Splatter Image [48] predict 3D Gaussians from image features, while MVSplat [9] encodes the feature matching information with cost volumes and achieves better geometry.
@@
 Another line of work like LGM [51], GRM [68], and GS-LRM [77] relies significantly on the training data and compute, discarding explicit feature matching cues and learning priors purely from data.
-This makes them expensive to train (e.g., GS-LRM [72] is trained with 64 A100 GPUs for 2 days), while our model can be trained in 20 hours with 8 GPUs.
-Moreover, our Gaussian splatting module, in return, enables pre-training depth model from large-scale unlabeled datasets without the need for ground truth depths.
+This makes them expensive to train (e.g., GS-LRM [77] is trained with 64 A100 GPUs for 2 days), while our model can be trained in 2 days with 4 GPUs.
+Moreover, our Gaussian splatting module enables pre-training depth model from large-scale multi-view posed datasets without ground truth depth supervision.
 Depth and Gaussian Splatting.
@@
 However, a crucial difference is that TranSplat uses monocular features to refine the coarse depth predicted from a cost volume, which makes it inherently vulnerable to the error in the coarse stage and, thus, suffers from the error propagation issue.
-In contrast, we avoid this by combining cost volume and monocular features early, which also enables our method to perform significantly better than TranSplat (Tab.
-5).
-There has been another line of work [10, 50] which applies an additional depth loss in the Gaussian splatting optimization process.
+In contrast, we avoid this by combining cost volume and monocular features early, which performs significantly better than TranSplat (Tab.
+6).
+Another line of work [12, 53] applies an additional depth loss to the Gaussian optimization process.
 We note that these two approaches (feed-forward vs.
@@
 3 DepthSplat
-Given $N$ input images $\{{\bm{I}}^{i}\}_{i=1}^{N},({\bm{I}}^{i}\in\mathbb{R}^{H\times W\times 3}$, where $H$ and $W$ are the image sizes) with corresponding projection matrices $\{\mathbf{P}_{i}\}_{i=1}^{N}$, ($\mathbf{P}_{i}\in\mathbb{R}^{3\times 4}$, computed from the intrinsic and extrinsic matrices), our goal is to predict dense per-pixel depth $\mathbf{D}_{i}\in\mathbb{R}^{H\times W}$ and per-pixel Gaussian parameters $\{(\bm{\mu}_{j},\alpha_{j},\bm{\Sigma}_{j},\bm{c}_{j})\}_{j=1}^{H\times W\times N}$ for each image, where $\bm{\mu}_{j}$, $\alpha_{j}$, $\bm{\Sigma}_{j}$ and $\bm{c}_{j}$ are the 3D Gaussian’s position, opacity, covariance, and color information.
+Given $N$ input images $\{{\bm{I}}^{i}\}_{i=1}^{N},({\bm{I}}^{i}\in\mathbb{R}^{H\times W\times 3}$, where $H$ and $W$ are the image sizes) with corresponding projection matrices $\{\mathbf{P}_{i}\}_{i=1}^{N}$, ($\mathbf{P}_{i}\in\mathbb{R}^{3\times 4}$, computed from the intrinsic and extrinsic matrices), our goal is to predict dense per-pixel depth $\mathbf{D}_{i}\in\mathbb{R}^{H\times W}$ and per-pixel Gaussian parameters $\{(\bm{\mu}_{j},\alpha_{j},\bm{\Sigma}_{j},\bm{c}_{j})\}_{j=1}^{H\times W\times N}$ for each image, where $\bm{\mu}_{j}$, $\alpha_{j}$, $\bm{\Sigma}_{j}$ and $\bm{c}_{j}$ are the 3D Gaussian’s position, opacity, covariance, and color.
 As shown in Fig.
-2, the core of our method is a multi-view depth model augmented with monocular depth features, where we obtain the position $\bm{\mu}_{j}$ of each Gaussian by unprojecting depth to 3D with camera parameters, and other Gaussian parameters are predicted by an additional lightweight head.
-More specifically, our depth model consists of two branches: one for modeling feature matching information using cost volumes, and another for extracting monocular features from a pre-trained monocular depth network.
+2, key to our method is a multi-view depth model augmented with monocular depth features, where we obtain the position $\bm{\mu}_{j}$ of each Gaussian by unprojecting depth with camera parameters, and other Gaussian parameters are predicted by an additional head.
+More specifically, our depth model consists of two branches: one for modeling feature matching using cost volumes, and another for extracting monocular features from a pre-trained monocular depth network.
 The cost volumes and monocular features are concatenated together for subsequent depth regression with a 2D U-Net and a softmax layer.
 For the depth task, we train our depth model with ground truth depth supervision.
-Our full model for novel view synthesis is trained with the photometric rendering loss, which can also be used as an unsupervised pre-training stage for the depth model.
-In the following, we introduce the individual components.
+Our full model for novel view synthesis is trained with the rendering loss, which can also be used as an unsupervised pre-training stage for the depth model.
+We introduce the individual components below.
 3.1 Multi-View Feature Matching
@@
 To exchange information across different views, we use a multi-view Swin Transformer [34, 64, 65] which contains six stacked self- and cross-attention layers to obtain multi-view-aware features $\{{\bm{F}_{i}}\}_{i=1}^{N}$, ${\bm{F}}^{i}\in\mathbb{R}^{\frac{H}{s}\times\frac{W}{s}\times C}$, where $C$ is the feature dimension.
-More specifically, cross-attention is performed between each reference view and other views.
 When more than two images ($N>2$) are given as input, we perform cross-attention between each reference view and its top-$2$ nearest neighboring views, which are selected based on their camera distances to the reference view.
-This makes the computation tractable while maintaining cross-view interactions.
+This makes computation tractable with many input views.
 Feature Matching.
@@
 Such a $2$-scale hierarchical architecture not only leads to improved efficiency since most computation is spent on low resolution, but also leads to better results thanks to the use of higher-resolution features [25].
-Similar feature fusion and depth regression procedure is used to get higher-resolution depth predictions, which are subsequently upsampled [38] to the full resolution.
+Similar feature fusion and depth regression procedure is used to get higher-resolution depth predictions, which are subsequently upsampled with a DPT head [41] to the full resolution.
 3.4 Gaussian Parameter Prediction
 For the task of 3D Gaussian splatting, we directly unproject the per-pixel depth maps to 3D with the camera parameters as the Gaussian centers $\bm{\mu}_{j}$.
-To predict other remaining Gaussian parameters $\alpha_{j}$ (opacity), $\bm{\Sigma}_{j}$ (covariance) and $\bm{c}_{j}$ (color), we use an additional lightweight 2D U-Net [39], where it takes the concatenated image, depth and feature information as input and outputs all the remaining Gaussian parameters.
+To predict other remaining Gaussian parameters $\alpha_{j}$ (opacity), $\bm{\Sigma}_{j}$ (covariance) and $\bm{c}_{j}$ (color), we use an additional DPT head [41], where it takes the concatenated image, depth and feature information as input and outputs all the remaining Gaussian parameters.
 With all the predicted 3D Gaussians, we can render novel views with the Gaussian splatting operation [29].
@@
 Training Details.
-For depth experiments, we mainly follow the setup of UniMatch [61] for fair comparisons on the ScanNet [12] dataset.
-More specifically, we train our depth model on 4 GH200 GPUs for 100K iterations with a total batch size of 32.
+For depth experiments on ScanNet [14], we train our depth model on 4×4$\times$4 × GH200 GPUs for 100K iterations with a total batch size of 32.
 For the ablation experiments on depth task, we mainly use synthetic datasets (TartanAir [58] and VKITTI2 [5]) for training since they provide high-quality ground truth depths.
@@
 We also report high-resolution ($512\times 960$) results on RealEstate10K and DL3DV datasets for qualitative evaluations.
-For experiments on the $256\times 256$ RealEstate10K dataset, we train our model on 8 GH200 GPUs for 150K iterations with a total batch size of 32, which takes about 20 hours.
-For experiments on the $256\times 448$ DL3DV [30] dataset, we evaluate on the official benchmark split with 140 scenes, while other remaining scenes are used for training.
-We fine-tune our RealEstate10K pre-trained model on 4 A100 GPUs for 100K iterations with a total batch size of 4, where the number of input views is randomly sampled from 2 to 6.
+For experiments on the $256\times 256$ RealEstate10K dataset, we train our model on 4×4$\times$4 × GH200 GPUs for 150K iterations with a total batch size of 32, which takes 1 day for the small model and 2 days for the large model.
+For experiments on the $256\times 448$ DL3DV [32] dataset, we fine-tune our RealEstate10K pre-trained model on 4×4$\times$4 × GH200 GPUs for 100K iterations with a total batch size of 4, where the number of input views is randomly sampled from 2 to 6.
 We evaluate the model’s performance on different number of input views (2, 4, 6).
@@
 3.3), where both larger monocular backbones and $2$-scale hierarchical models lead to consistently improved performance for both depth and view synthesis tasks.
-Our model parameters, inference time, and memory consumption are provided in Tab.
-A.1 of the appendix.
-Monocular Multi-View Depth (TartanAir) 3DGS (RealEstate10K)
+The inference time is measured on a single A100 GPU for two input views at $256\times 256$ resolutions.
+Mono Multi-View Depth (TartanAir) 3DGS (RealEstate10K) Param (M) Time (s)
 Abs Rel $\downarrow$ $\delta_{1}\uparrow$ PSNR $\uparrow$ SSIM $\uparrow$ LPIPS $\downarrow$
-ViT-S $1$-scale 8.46 93.02 26.76 0.877 0.123
-ViT-B $1$-scale 6.94 94.46 27.09 0.881 0.119
-ViT-L $1$-scale 6.07 95.52 27.34 0.885 0.118
-ViT-S $2$-scale 7.01 94.56 26.96 0.880 0.122
-ViT-B $2$-scale 6.22 95.31 27.27 0.885 0.120
-ViT-L $2$-scale 5.57 96.07 27.44 0.887 0.119
+ViT-S $1$-scale 8.46 93.02 26.84 0.878 0.122 38 0.050
+ViT-B $1$-scale 6.94 94.46 27.06 0.882 0.119 115 0.053
+ViT-L $1$-scale 6.07 95.52 27.23 0.885 0.116 347 0.061
+ViT-S $2$-scale 7.01 94.56 27.12 0.884 0.119 41 0.062
+ViT-B $2$-scale 6.22 95.31 27.34 0.887 0.116 120 0.070
+ViT-L $2$-scale 5.57 96.07 27.47 0.889 0.114 354 0.079
 Table 2: Ablations.
@@
 Abs Rel $\downarrow$ $\delta_{1}\uparrow$ PSNR $\uparrow$ SSIM $\uparrow$ LPIPS $\downarrow$
-full 8.46 93.02 26.76 0.877 0.123
-w/o mono feature 12.25 88.00 26.27 0.866 0.130
-w/o cost volume 11.34 90.02 23.09 0.761 0.187
-single branch 11.26 90.84 25.99 0.858 0.134
-ConvNeXt-T 10.50 91.13 26.28 0.867 0.130
-Midas 9.53 91.61 26.40 0.869 0.129
-DINO V2 8.93 92.49 26.68 0.874 0.125
-Depth Anything V1 8.38 93.23 26.70 0.875 0.125
-Depth Anything V2 8.46 93.02 26.76 0.877 0.123
+Full 8.46 93.02 26.84 0.878 0.122
+w/o mono feature 12.25 88.00 26.04 0.864 0.134
+w/o cost volume 11.34 90.02 23.24 0.766 0.184
+ConvNeXt-T 10.50 91.13 26.11 0.865 0.134
+Midas 9.53 91.61 26.46 0.872 0.128
+DINO V2 8.93 92.49 26.76 0.877 0.123
+Depth Anything V1 8.38 93.23 26.76 0.877 0.124
+Depth Anything V2 8.46 93.02 26.84 0.878 0.122
 Figure 3: Effect of monocular features for depth on ScanNet.
@@
 1 demonstrate that both larger monocular backbones and $2$-scale hierarchical models lead to consistently improved performance for both tasks.
+We name three different model sizes with “Small": “ViT-S monocular, 1-scale multi-view" , “Base": “ViT-B monocular, 2-scale multi-view", and “Large": “ViT-L monocular, 2-scale multi-view".
 From Tab.
 1, we can also observe that better depth network architecture leads to improved view synthesis.
-In Tab.
-B.1 of the appendix, we conduct additional experiments to study the effect of different initializations for the depth network to the view synthesis performance.
+In the appendix, we conduct additional experiments to study the effect of different initializations for the depth network to the view synthesis performance.
 Our results show that a better depth model initialization also contributes to improved rendering quality.
@@
 This indicates that obtaining scale- and multi-view consistent predictions with a pure monocular depth backbone is challenging, which constrains achieving high-quality results for 3D Gaussian reconstruction.
-Fusion Strategy.
-We compare with alternative strategies for fusing the monocular features for multi-view depth estimation.
-In particular, we compare with MVSFormer [6] which constructs the cost volume with monocular features.
-More specifically, we replace our multi-view feature extractor with a weight-sharing ViT model and use the ViT features to build the cost volume as done in MVSFormer.
-This leads to a single-branch architecture (single branch in Tab.
-2), unlike our two-branch design where the monocular features are not used to build the cost volume.
-We can observe that our fusion strategy performs significantly better than the single-branch design, potentially because our two-branch design disentangles feature matching and obtaining monocular priors, which makes the learning task easier.
+Table 3: Monocular fusion strategy. Our simple concatenation performs surprisingly good compared to other alternatives.
+Method Abs Rel $\downarrow$ $\delta_{1}\uparrow$
+Cost volume from ViT 11.26 90.84
+Explicit scale align 9.13 90.21
+Attention fusion 8.40 92.95
+Concatenation (Ours) 8.46 93.02
 Table 4: Unsupervised Depth Pre-Training with Gaussian Splatting.
@@
 Method Abs Rel $\downarrow$ RMSE $\downarrow$ RMSElog $\downarrow$
-DeMoN [51] 0.231 0.761 0.289
-BA-Net [47] 0.161 0.346 0.214
-DeepV2D [49] 0.057 0.168 0.080
-NeuralRecon [44] 0.047 0.164 0.093
-DRO [24] 0.053 0.168 0.081
-UniMatch [61] 0.059 0.179 0.082
-DepthSplat 0.044 0.119 0.059
+DeMoN [54] 23.1 0.761 0.289
+BA-Net [50] 16.1 0.346 0.214
+DeepV2D [52] 5.7 0.168 0.080
+NeuralRecon [47] 4.7 0.164 0.093
+DRO [26] 5.3 0.168 0.081
+UniMatch [65] 5.9 0.179 0.082
+DepthSplat (w/o GS pre-train) 4.5 0.125 0.061
+DepthSplat (w/ GS pre-train) 3.8 0.114 0.055
 Different Monocular Features.
@@
 The pre-trained Depth Anything V2 [70] monocular features achieve the best view synthesis results.
-4.3 Unsupervised Depth Pre-Training with Gaussian Splatting
-By connecting Gaussian splatting and depth, our DepthSplat provides a way to pre-train the depth model in a fully unsupervised manner.
-In particular, we first train our full model on the large-scale unlabeled RealEstate10K dataset (contains $\sim$ 67K Youtube videos) with only the Gaussian splatting rendering loss (Eq.
-2), without any direct supervision on the depth predictions.
-After pre-training, we take the pre-trained depth model and further fine-tune it to the depth task on the mixed TartanAir [55] and VKITTI2 [5] datasets with ground truth depth supervision.
-In Tab.
-3, we evaluate the performance on both in-domain TartanAir test set and the zero-shot generalization on unseen ScanNet and KITTI datasets.
-We can observe that the benefit of pre-training is especially significant on the challenging datasets like TartanAir and KITTI.
-The results align well with the observation in [19] which suggests that the pre-training acts as regularization and guides the learning towards better minima and enables better generalization to out-of-distribution datasets.
-The visual comparison results are provided in Fig.
-C.1 of the appendix.
-Given the increasing popularity of view synthesis [74, 57] and multi-view generative models [42, 3], new multi-view datasets [30] and models [52] are gradually introduced, our approach provides a way to pre-train depth models on large-scale unlabeled multi-view image datasets.
+Monocular Fusion Strategy.
+We compare with alternative strategies for fusing the monocular features for multi-view depth estimation in Tab.
+3.
+In particular, we first compare with MVSFormer [6] that constructs the cost volume with monocular ViT features, which leads to a single-branch architecture.
+We can observe that our fusion strategy performs significantly better than the single-branch design, potentially because our two-branch design disentangles feature matching and monocular priors, which makes the learning task easier.
+We also compare with fusing the monocular depth maps (instead of monocular features in our model) by first aligning the monocular depth scale using the multi-view depth and then fusing the depth maps together with an additional network.
+Again, our simple concatenation performs better.
+We further compare our concatenation with the attention-based adaptive fusion and observe very similar results.
+Thus we opt to the simpler concatenation design.
 Table 6: Two-view 3DGS on RealEstate10K. Our DepthSplat achieves the best performance.
@@
 TranSplat [76] 26.69 0.875 0.125
-DepthSplat 27.44 0.887 0.119
-Figure 5: Visual synthesis on RealEstate10K. Our DepthSplat performs significantly better than pixelSplat [7] and MVSplat [8] in challenging regions.
-Figure 6: Depth results on unseen samples. Our depth model generalizes robustly to different scene types.
+DepthSplat 27.47 0.889 0.114
+Figure 5: Novel view synthesis on RealEstate10K. Our DepthSplat performs significantly better than pixelSplat [7] and MVSplat [9] in challenging regions.
+4.3 Unsupervised Depth Pre-Training with Gaussian Splatting
+By connecting Gaussian splatting and depth, our DepthSplat provides a way to pre-train the depth model in a fully unsupervised manner.
+In particular, we first train our full model on the large-scale multi-view posed RealEstate10K dataset (contains $\sim$ 67K Youtube videos) with only the Gaussian splatting rendering loss (Eq.
+2), without any direct supervision on the depth predictions.
+After pre-training, we take the pre-trained depth model and further fine-tune it to the depth task on the mixed TartanAir [58] and VKITTI2 [5] datasets with ground truth depth supervision.
+In Tab.
+4, we evaluate the performance on both in-domain TartanAir test set and the zero-shot generalization on unseen ScanNet and KITTI datasets.
+We can observe that the benefit of pre-training is especially significant on the challenging datasets like TartanAir and KITTI.
+The results align well with the observation in [21] which suggests that the pre-training acts as regularization and guides the learning towards better minima and enables better generalization to out-of-distribution datasets.
+The visual comparison results are provided in the appendix.
+Given the increasing popularity of view synthesis [81, 61] and multi-view generative models [45, 3], new multi-view datasets [32] and models [55] are gradually introduced, our approach provides a way to pre-train depth models on large-scale multi-view posed image datasets.
 4.4 Benchmark Comparisons
@@
 5, where our method significantly improves the performance on challenging scenarios like texture-less regions and occlusions.
+For the depth results on ScanNet, we report our base models (with ViT-B and two-scale feature matching) trained with and without Gaussian splatting pre-training on the mixed RealEstate10K and DL3DV datasets.
+Our model significantly outperforms previous methods and can benefit from additional Gaussian splatting pre-training.
 Table 7: Comparisons on DL3DV. Our DepthSplat not only consistently outperforms MVSplat on different number of input views, but also scales more efficiently to more input views.
@@
 MVSplat [9] 2 17.54 0.529 0.402 0.072
-DepthSplat 19.05 0.610 0.313 0.101
+DepthSplat 19.31 0.615 0.310 0.083
 MVSplat [9] 4 21.63 0.721 0.233 0.146
-DepthSplat 22.82 0.766 0.188 0.124
+DepthSplat 23.12 0.780 0.178 0.107
 MVSplat [9] 6 22.93 0.775 0.193 0.263
-DepthSplat 23.83 0.808 0.158 0.161
+DepthSplat 24.19 0.823 0.147 0.132
 Figure 6: Visual comparisons on DL3DV. Our DepthSplat performs significantly better than MVSplat [9] on regions that hard to match (e.g., repeated patterns).
@@
 We fine-tune MVSplat and our RealEstate10K pre-trained models on DL3DV training scenes and report the results on the benchmark test set in Tab.
-6.
+7.
 Our DepthSplat consistently outperforms MVSplat in all metrics, and our improvements are more significant with fewer input views, which indicates that our model is more robust to sparse views.
 Visual comparisons with MVSplat on the DL3DV dataset are shown in Fig.
-7, where MVSplat’s depth quality lags behind our DepthSplat due to matching failure, which leads to blurry and distorted view synthesis results.
-We show more visual comparison results in Fig.
-C.3 of the appendix.
+6, where MVSplat’s depth quality lags behind our DepthSplat due to matching failure, which leads to blurry and distorted view synthesis results.
+We show more visual comparison results in the appendix.
 It is also worth noting that our method scales more efficiently to more input views thanks to our lightweight local feature matching approach (Sec.
 3.1), which is unlike the expensive global pair-wise matching used in MVSplat.
-Table 7: Cross-dataset generalization. When generalizing to DL3DV with RealEstate10K pre-trained models, our DepthSplat is consistently better than MVSplat for both small and large baselines.
-Method Small Baseline Large Baseline
+Table 8: Cross-dataset generalization.
+When generalizing to DL3DV and ACID with RealEstate10K pre-trained models, our DepthSplat is consistently better than MVSplat, especially on the more challenging DL3DV dataset.
+Method DL3DV ACID
 PSNR $\uparrow$ SSIM $\uparrow$ LPIPS $\downarrow$ PSNR $\uparrow$ SSIM $\uparrow$ LPIPS $\downarrow$
-MVSplat [8] 19.19 0.608 0.304 16.26 0.459 0.438
-DepthSplat 20.29 0.672 0.259 17.50 0.536 0.376
+MVSplat [9] 24.14 0.808 0.149 28.15 0.841 0.147
+DepthSplat 27.66 0.895 0.096 28.37 0.847 0.141
 Cross-Dataset Generalization.
-We further evaluate the cross-dataset generalization capability of our model by directly testing on DL3DV with RealEstate10K pre-trained models.
+We evaluate the cross-dataset generalization capability by directly testing on DL3DV [32] and ACID [33] datasets with RealEstate10K [82] pre-trained models.
 The results are shown in Tab.
-7, where our DepthSplat consistently outperforms MVSplat for both small and large baseline inputs (visual comparisons are shown in Fig.
-C.2 of the appendix).
-We also show some visual results of our depth model on unseen samples in Fig.
-6, where our model generalizes robustly to different scene types.
+8, where our DepthSplat consistently outperforms MVSplat on both datasets, especially on the more challenging DL3DV dataset which features complex scene structures.
+The visual comparison results are shown in the appendix.
 High-Resolution Results.
-In addition to the comparison results on the $256\times 256$ resolution RealEstate10K and $256\times 448$ DL3DV dataets, we show more high-resolution ($512\times 960$) qualitative results in Fig.
-C.4 and Fig.
-C.5 of the appendix.
-We also invite the readers to our project page haofeixu.github.io/depthsplat for high-resolution ($512\times 960$) video results on different number of input views (6 and 12), where our DepthSplat is able to reconstruct larger-scale or 360 scenes from more input views.
+In addition to the comparison results on the $256\times 256$ resolution RealEstate10K and $256\times 448$ resolution DL3DV dataets, we show more high-resolution ($512\times 960$) qualitative results in the appendix.
+We also invite the readers to our project page haofeixu.github.io/depthsplat for high-resolution ($512\times 960$) video results on different number of input views (6 and 12), where our DepthSplat is able to reconstruct larger-scale or $360^{\circ}$ scenes from more input views.
 5 Conclusion
 In this paper, we introduce DepthSplat, a new approach to connecting Gaussian splatting and depth to achieve state-of-the-art results on ScanNet, RealEstate10K and DL3DV datasets for both depth and view synthesis tasks.
-We also show that our model enables unsupervised pre-training depth with Gaussian splatting rendering loss, providing a way to leverage large-scale unlabeled multi-view image datasets for training more robust and generalizable multi-view depth models.
-Our current model requires camera pose information as input along with the multi-view images, removing this requirement would be exciting future work.
-6 Acknowledgement
-We thank Yuedong Chen for his generous help with the DL3DV dataset. Andreas Geiger was supported by the ERC Starting Grant LEGO-3D (850533) and the DFG EXC number 2064/1 - project number 390727645.
+We show that our model enables unsupervised depth pre-training with Gaussian splatting rendering loss, providing a way to leverage large-scale posed multi-view image datasets for training more robust and generalizable multi-view depth models.
+Limitations.
+Our current model requires camera poses as input along with the multi-view images, which might be challenging to obtain when the input views are extremely sparse.
+An interesting future direction is to remove this requiement by exploring pose-free models [72, 80].
+In addition, our model predicts pixel-aligned Gaussians, where the number of Gaussians will be significantly large when processing a large number of input views [60, 79].
+Further improving the geometry representation and the scalability to many input views would be exciting future work.
+Acknowledgement
+We thank Yuedong Chen for his generous help with the DL3DV dataset.
+Andreas Geiger was supported by the ERC Starting Grant LEGO-3D (850533) and the DFG EXC number 2064/1 - project number 390727645.
+This work was supported as part of the Swiss AI Initiative by a grant from the Swiss National Supercomputing Centre (CSCS) under project ID a03 on Alps.
+H.
+Blum is also with the Lamarr Institute and D.
+Barath is also with HUN-REN SZTAKI.
 Appendix
-A Model Efficiency
-We report the number of parameters, inference time and memory consumption in Tab.
-A.1 for different model sizes.
-In particular, the “small” model refers to “ViT-S monocular, 1-scale multi-view” in Tab.
-1 of the main paper, and the “base” model refers to “ViT-B monocular, 2-scale multi-view”, and the “large” model refers to “ViT-L monocular, 2-scale multi-view”.
-We report the inference time (measured on an A100 GPU) and memory for 2 input views at $256\times 256$ and $512\times 512$ resolutions.
-Our DepthSplat can support different speed-accuracy trade-offs.
-Table A.1: Model efficiency.
-Model Size Resolution Params (M) Time (s) Memory (GB)
-Small $256\times 256$ 37 0.065 1.4
-Base 117 0.081 2.1
-Large 360 0.105 3.5
-Small $512\times 512$ 37 0.091 3.1
-Base 117 0.155 5.4
-Large 360 0.313 8.4
-B Depth Pre-Training for Gaussian Splatting
-In Sec.
-4.1 of the main paper, we have shown that better depth architectures lead to improved view synthesis with Gaussian splatting.
+A Depth Pre-Training for Gaussian Splatting
+In the main paper, we have shown that better depth architectures lead to improved view synthesis with Gaussian splatting.
 We further study the effect of different weight initializations for the depth model.
@@
 We can observe from Tab.
-B.1 that better depth initialization also leads to improved view synthesis results.
-Table B.1: Depth pre-training for Gaussian splatting.
+A.1 that better depth initialization also leads to improved view synthesis results.
+Table A.1: Depth pre-training for Gaussian splatting.
 We compare different weight initializations for the depth model when training our full DepthSplat model for view synthesis.
@@
 full depth model 26.81 0.878 0.1225
-C More Visual Results
-C.1 Unsupervised Depth Pre-Training with Gaussian Splatting
+B More Visual Results
+B.1 Unsupervised Depth Pre-Training with Gaussian Splatting
 In Fig.
-C.1, we show the comparisons of depth estimation results with and without Gaussian splatting pre-training.
+B.1, we show the comparisons of depth estimation results with and without Gaussian splatting pre-training.
 We observe that pre-training leads to better results in texture-less regions.
 We hypothesize that pre-training provides regularizations (as also observed in  [21]) to the challenging scenarios and accordingly leads to improved performance.
-Figure C.1: Effect of unsupervised depth pre-training. We observe that the unsupervised pre-training leads to improved performance for texture-less regions.
-C.2 Cross-Dataset Generalization
+Figure B.1: Effect of unsupervised depth pre-training. We observe that the unsupervised pre-training leads to improved performance for texture-less regions.
+B.2 Cross-Dataset Generalization
 In Fig.
-C.2, we show the cross-dataset generalization results on the DL3DV [30] dataset, which are obtained with the RealEstate10K [75] trained models.
+B.2, we show the cross-dataset generalization results on the DL3DV [32] dataset, which are obtained with the RealEstate10K [82] pre-trained models.
 Our DepthSplat generalizes more robustly than MVSplat [9] on unseen scenes.
-Figure C.2: Generalization from RealEstate10K to DL3DV.
-C.3 Visual Comparisons on DL3DV
-In Fig. C.3, we compare the visual synthesis results from 4 input views with MVSplat [8]. Our DepthSplat better preserves the scene structures.
-Figure C.3: View synthesis from 4 input views on DL3DV.
-C.4 High-Resolution Results
+Figure B.2: Generalization from RealEstate10K to DL3DV.
+B.3 Visual Comparisons on DL3DV
+In Fig. B.3, we compare the visual synthesis results from 4 input views with MVSplat [9]. Our DepthSplat better preserves the scene structures.
+Figure B.3: View synthesis from 4 input views on DL3DV.
+B.4 High-Resolution Results
 In Fig.
-C.4 and Fig.
-C.5, we show the view synthesis and multi-view depth estimation results at $512\times 960$ resolutions.
+B.4 and Fig.
+B.5, we show the view synthesis and multi-view depth estimation results at $512\times 960$ resolutions.
 Note that MVSplat [9] runs out-of-memory on such high resolutions with 6 or 12 input views, while our DepthSplat significantly improves the efficiency with two technical components.
 First, our lowest feature resolution is $1/8$ of the image resolution, while MVSplat uses $1/4$.
-Second, we use local cross-view attentions (detailed in Sec.
-3.1 of the main paper) unlike the pair-wise global attentions in MVSplat.
-Figure C.4: View Synthesis at $512\times 960$ resolutions from 6 input views.
-Figure C.5: Depth predictions on DL3DV with 12 input views. The image resolutions are $512\times 960$.
-D More Implementation Details
-We provide more implementation details on high-resolution experiments.
+Second, we use local cross-view attentions unlike the pair-wise global attentions in MVSplat.
+Figure B.4: View synthesis at $512\times 960$ resolutions from 6 input views.
+Figure B.5: Depth predictions on DL3DV with 12 input views. The image resolutions are $512\times 960$.
+C More Implementation Details
+We provide more implementation details on the high-resolution experiments.
 For high-resolution experiments, we choose our small model which contains a ViT-S monocular branch and a single-scale multi-view branch.
-We first train our model on RealEstate10K at $448\times 768$ resolutions for 200K iterations with a batch size of 8, and then we fine-tune the model on DL3DV at $448\times 768$ resolutions for 200K iterations with a batch size of 8.
-We use random number of input views (2 to 6) for training.
-Despite that our model is trained at $448\times 768$ resolutions with at most 6 input views, we observe our model can generalize reasonably well on higher resolutions (e.g., $512\times 960$) and more input views (e.g., 12), as we show in our project page: haofeixu.github.io/depthsplat.
+We first train our model on RealEstate10K with two input views at $256\times 448$ resolutions for 150K iterations, where the total batch size is 256.
+We fine-tune the model on the mixed RealEstate10K and DL3DV datasets at $448\times 768$ resolutions for 200K iterations with a total batch size of 64, where the number of input views is randomly sampled from 4 to 10.
+We use this model to predict results on high resolutions (e.g., $512\times 960$) and different numbers of input views (e.g., 6 and 12), as we show in our project page: haofeixu.github.io/depthsplat.